### v4

Abstract
We propose and analyze a randomization scheme for a general class of impulse control problems. The solution to this randomized problem is characterized as the fixed point of a compound operator which consists of a regularized nonlocal operator and a regularized stopping operator. This approach allows us to derive a semi-linear Hamilton-Jacobi-Bellman (HJB) equation. Through an equivalent randomization scheme with a Poisson compound measure, we establish a verification theorem that implies the uniqueness of the solution. Via an iterative approach, we prove the existence of the solution. The existence–and–uniqueness result ensures the randomized problem is well-defined. We then demonstrate that our randomized impulse control problem converges to its classical counterpart as the randomization parameter $\bm{\lambda}$ vanishes. This convergence, combined with the value function’s $\mathcal{C}^{2,\alpha}_{loc}$ regularity, confirms our framework provides a robust approximation and a foundation for developing learning algorithms. Under this framework, we propose an offline reinforcement learning (RL) algorithm. Its policy improvement step is naturally derived from the iterative approach from the existence proof, which enjoys a geometric convergence rate. We implement a model-free version of the algorithm and numerically demonstrate its effectiveness using a widely-studied example. The results show that our RL algorithm can learn the randomized solution, which accurately approximates its classical counterpart. A sensitivity analysis with respect to the volatility parameter $\sigma$ in the state process effectively demonstrates the exploration–exploitation tradeoff.
Keywords: impulse control, nonlocal operator, optimal stopping, entropy-regularized randomization, reinforcement learning
1 Introduction
Impulse control models systems in which actions are infrequent but high-impact: a controller waits and then intervenes at selected times to shift the state by a chosen jump size. The policy usually takes the form of $\nu=\{(\tau_{n},\xi_{n})\}_{n\geq 1}$, where $\{\tau_{n}\}_{n\geq 1}$ denotes a sequence of stopping times indicating the timing of inventions, and $\{\xi_{n}\}_{n\geq 1}$ denotes the associated strategic jumps. Under a given policy $\nu$, the controlled state dynamics is given by
$\mathrm{d}X_{t}=b(X_{t-})\mathrm{d}t+\sigma(X_{t-})\mathrm{d}W_{t}+\sum_{n\geq 1}\xi_{n}\delta(t-\tau_{n}).$
The objective is to minimize the discounted total cost,
$\psi(x)=\inf_{\nu}J(x;\nu)=\inf_{\nu}\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t-})\mathrm{d}t+\sum_{n\geq 1}e^{-r\tau_{n}}l(\xi_{n})\Big|X_{0-}=x\right],$
with a discount rate $r>0$, a running cost $f$, and an intervention cost $l$. Compared to regular control, impulse control is a more versatile mathematical framework as it incorporates discontinuities into the state processes. It also generalizes optimal stopping or the singular control problems, as impulse interventions can incur a nonzero fixed cost HT1983 . As a result, various problems in mathematical finance, engineering and management can be modeled as impulse control problems, including cash management constantinides1978existence , inventory controls Harrison1983 ; Sulem1986 , transaction cost in portfolio analysis Eastham1988 ; Morton1995 ; Korn1998 ; Korn1999 ; Bielecki2000 ; Oksendal2002 , insurance model Jeanblanc-Picque1995 ; Candenillas2006 , liquidity risk LyVath2007 , exchange rates Jeanblanc-Picque1993 ; Mundaca1998 ; Bertola2016 , real options Triantis1990 ; Mauer1994 ; BensoussanRoignant , energy market basei2019optimal , and so on.
In contrast to the popularity as a modeling tool, impulse control problems face technical difficulties in deriving and analyzing the optimal strategies. In general, the analysis of impulse control problems is often through the associated Hamilton-Jacobi-Bellman (HJB) equation of the quasi-variational inequality (QVI) type,
$\min\left\{\frac{\sigma^{2}}{2}\psi^{\prime\prime}+b\psi^{\prime}-r\psi+f,\mathcal{M}\psi-\psi\right\}=0,$
where $\mathcal{M}$ is a nonlocal operator such that $\mathcal{M}\psi(x)=\inf_{\xi\in\mathbb{R}}\{\psi(x+\xi)+l(\xi)\}$; see, for instance, bensoussan1984impulse ; oksendal2007applied for more details on the QVIs. Due to the very presence of this nonlocal operator and correspondingly the discontinuous nature of the controlled process, it takes delicate treatments of the value function, especially its regularity conditions, so that popular methods such as the smooth-fit principle can be applied; see for instance guo2009smooth ; davis2010impulse ; bayraktar2013impulse . One of the best known strategies would be the threshold-type $(d,D,U,u)$ derived for the one-dimensional cash management model in constantinides1978existence : the controller intervenes only if the state process exits the continuation region $(d,u)$, and in this case, the state is immediately adjusted to the closer threshold between $D$ and $U$; see also alvarez2004class for more explicitly solvable cases. It is particularly challenging to explicitly derive the value function and the corresponding optimal strategy under a generic impulse control framework.
To numerically solve impulse control problems instead, several numerical methods for the corresponding HJB-QVIs have been developed. Many of these methods are based on iterative partial differential equation (PDE) frameworks. The most common route is through penalization combined with policy/Howard iteration: one solves a sequence of penalized HJB equations and iterates to a fixed point (see, e.g., bensoussan1984impulse ; reisinger2019penalty ; reisinger2020error ). Other widely used approaches include direct-control or semi-Lagrangian discretizations with policy iteration and matrix-monotonicity arguments azimzadeh2016weakly ; azimzadeh2018convergence , finite-difference (FD)/finite-element (FE) schemes that provide $L^{\infty}$-type error bounds for (non)coercive QVIs boulbrachene1998noncoercive ; boulbrachene2005pointwise , implicit time-stepping for finite-horizon formulations ieda2015implicit , and probabilistic representations via backward stochastic differential equations (BSDEs) that enable Monte-Carlo solvers for certain classes kharroubi2010backward . While effective for low dimensional state space, these methods face serious limitations in modern applications: grid-based PDE schemes suffer from the curse of dimensionality and delicate monotonicity/stencil requirements; penalty formulations introduce stiffness and parameter tuning (often degrading accuracy near free boundaries); FD/FE error analyses can hinge on regularity or coercivity not guaranteed in practice; and BSDE solvers may require nested simulation and restrictive jump structures.
As an alternative way to bypass the aforementioned obstacles in traditional numerical methods, reinforcement learning (RL) has been widely studied in recent years to solve optimal stochastic control problems under the continuous setting. The essence of RL is to learn about the possibly unknown environment from either the historical trajectory data (offline) or the interactions (online), and then to solve a decision making problem according to some objective function. Therefore, it is crucial to balance between learning and decision-making, also known as the “exploration–exploitation” tradeoff. To encourage exploration, a randomized optimal control framework is necessary. For regular control problems, a common randomization scheme is through an entropy regularization over the policy; see, for instance, wang2020reinforcement . Due to the newly-introduced Shannon’s entropy term, the controller is encouraged to explore the entire trajectory space instead of only exploiting the optimal one under the deterministic setting. In wang2020reinforcement , a thorough analysis is conducted on a linear-quadratic problem, where the exploration–exploitation tradeoff is full characterized through the optimal randomized feedback strategy as a Gaussian distribution. The analysis of the entropy-regularized framework under a general setting is conducted in tang2022exploratory . Under this framework, model-free updates have been developed in continuous-time temporal-difference learning and $q$-learning jia2022policy ; jia2023q , where the control randomization provides a policy-gradient estimator and has a variance reduction effect in the meantime.
To design learning algorithms for optimal stopping problems, a proper randomized framework essentially increases the frequency of receiving the otherwise extremely sparse reward signals associated with the stopping. One way to achieve that is to randomize the survival process $\{\mathbbm{1}\{\tau>t\}\}_{t\geq 0}$ for any given stopping time $\tau$. For instance, in dong2024randomized , the survival process is modeled as $\{e^{-\int_{0}^{t}\pi_{s}ds}\}_{t\geq 0}$ that is controlled by the nonnegative survival rate $\pi=\{\pi_{t}\}_{t\geq 0}$. With a running normalized entropy regularization $\mathcal{R}(\pi_{t})$, the randomized optimal stopping problem is solved via analyzing the associated semi-linear HJB equation. A corresponding offline RL algorithm is proposed. In dianetti2024exploratory , the survival process is modeled as $\{1-\xi_{t}\}_{t\geq 0}$ where $\xi=\{\xi_{t}\}_{t\geq 0}$ is a nondecreasing singular control process valued on $[0,1]$. With a cumulative relative entropy regularization on $\xi$, the randomized optimal stopping problem is formulated as a singular control problem whose state space is augmented with the current survival probability $y\in[0,1]$; its solution can be derived through its obstacle-type HJB equation. Correspondingly, two versions of RL algorithms, one model-based and the other model-free, are proposed to learn the decision boundary characterizing the optimal bang-bang control. In both dong2024randomized and dianetti2024exploratory , the respective randomized optimal stopping frameworks via the survival process are shown to converge to their classical counterparts as the parameter $\lambda$ modulating the randomization decreases to 0.
Extending these randomization ideas to impulse control problems is nontrivial, as control policy concerns with not just one but a sequence of stopping times together with the associated jumps. In denkert2025control , a special case of impulse control problem is considered where the policy $\nu$ is modeled as a marked point process with an associated (uncontrolled) Poisson random measure; see kharroubi2010backward for more details. Instead of randomizing the survival processes for each stopping time, the randomization in denkert2025control is through a controlled intensity process that tilts the probability measure. Under the intensity policy, the policy gradient is derived and a corresponding actor-critic RL algorithm is then proposed.
Our Work
In this work, we seek for a randomization scheme for a general class of impulse control problems whose control policies are beyond the form of marked point processes. The idea is inspired by the fixed-point characterization of the impulse control value function $\psi$ under the operator $\mathcal{T}=T\circ\mathcal{M}$, such that for any proper test function $\phi$,
$\mathcal{T}\phi(x)=T[\mathcal{M}\phi](x)=\inf_{\tau}\mathbb{E}\left[\int_{0}^{\tau}e^{-rt}f(Y_{t-})dt+e^{-r\tau}\mathcal{M}\phi(Y_{\tau-})\biggl|Y_{0-}=x\right],\ \ \forall x\in\mathbb{R},$
subject to $dY_{t}=b(Y_{t})dt+\sigma(Y_{t})dW_{t}$; see bensoussan1984impulse ; guo2009smooth . We propose a two-fold randomization scheme for impulse control problem. First, we randomize the nonlocal operator into $\mathcal{M}^{\lambda}$ for some $\lambda>0$ such that for any suitable test function $\phi$,
$\mathcal{M}^{\lambda}\phi(x)=\inf_{\mu,\mu\ll\Phi_{x}}\mathbb{E}_{\xi\sim\mu}\left[\phi(x+\xi)+l(\xi)\right]+\lambda D_{{\rm KL}}\left(\mu\|\Phi_{x}\right),\ \ \forall x\in\mathbb{R},$
where $\Phi_{x}$ denotes the Gaussian distribution $\mathcal{N}(-x,1)$, and $D_{{\rm KL}}\left(\mu\|\Phi_{x}\right)$ denotes the Kullback-Leibler divergence between $\mu$ and $\Phi_{x}$; here, $D_{{\rm KL}}\left(\mu\|\Phi_{x}\right)$ effectively acts as the entropy-regularizer that encourages the exploration of jumps. Then, we randomize the optimal stopping operator into $T^{\lambda}$ for some $\lambda>0$, such that for any suitable test function $\phi$,
$T^{\lambda}\phi(x)=\inf_{\pi_{t}\geq 0}\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}r+\pi_{s}ds}(f(Y_{t})-\lambda\mathcal{R}(\pi_{t})+\pi_{t}\phi(Y_{t}))dt\biggl|Y_{0-}=x\right],\ \ \forall x\in\mathbb{R},$
where $\mathcal{R}(\pi_{t})$ is the normalized entropy as in dong2024randomized to encourage the exploration of intervention times. Given any $\bm{\lambda}=(\lambda_{1},\lambda_{2})\in(0,+\infty)^{2}$, the value function $\psi^{\bm{\lambda}}$ of the randomized impulse control problem is the fixed point for the compound operator $\mathcal{T}^{\bm{\lambda}}=T^{\lambda_{1}}\circ\mathcal{M}^{\lambda_{2}}$; the intensity-jump distribution pair $(\pi^{*},\mu^{*})$ denotes the corresponding optimal optimal policy if $\psi^{\bm{\lambda}}$ is attained at $(\pi^{*},\mu^{*})$.
Based on the randomized framework introduced above, we first derive the associated HJB equation for the value function $\psi^{\bm{\lambda}}$. Then, using an equivalent randomization method via a Poisson compound measure, we provide a verification theorem. This verification theorem guarantees the uniqueness of the solution to the randomized impulse control problem. To establish the existence of the solution, we propose a constructive proof via an iterative approach. We then show that the iterates of value functions admit a pointwise limit, which is exactly a fixed point for the compound operator $\mathcal{T}^{\bm{\lambda}}$; indeed, the randomized impulse control value function $\psi^{\bm{\lambda}}$ is shown to be a $C^{2,\alpha}_{\mathrm{loc}}$ classical solution to the corresponding HJB equation, and it can be attained at a unique optimal randomized policy $(\pi^{*},\mu^{*})$. With some additional boundedness assumptions, we establish the convergence of $\psi^{\bm{\lambda}}$ to the corresponding classical impulse control value function $\psi$ as $\bm{\lambda}$ approaches $(0,0)$ upon taking a proper subsequence.
The iterative approach in the proof of existence for $\psi^{\bm{\lambda}}$ naturally gives rise to the corresponding RL algorithm, where the iterates correspond to the policy improvement steps whose convergence to the randomized value function is guaranteed by this existence proof. Under the same set of bounded assumptions, we can further derive a geometric convergence rate. We then design a temporal–difference (TD)-based algorithm. It provides an offline model-free learning procedure to solve the randomized impulse control problem. We implement and evaluate this RL algorithm on a classical linear model that has been widely explored in the impulse–control literature; see constantinides1978existence ; jeanblanc1993impulse ; guo2009smooth , for instance. The experiments empirically verify the convergence of the randomized problem to the classical one as $(\lambda_{1},\lambda_{2})$ approaches $(0,0)$. We also provide a sensitivity analysis with respect to the volatility parameter $\sigma$ in the linear model under the randomized setting. The exploration–exploitation trade off is demonstrated through the impact of $\sigma$ on value functions, intervention intensities as well as jump distributions.
The paper is organized as follows: In Section 2, we introduce a randomization framework for the impulse-control problem using some newly-defined entropy-regularized operators, and state the definition of the solution to this randomized problem. In Section 3, we discuss some preliminary properties for the aforementioned randomized operators. We present the HJB characterization as well as a verification theorem for the randomized impulse control problem in Section 4. This verification theorem guarantees the uniqueness of the solution to this randomized solution. In Section 5, we propose an iterative approach to prove the existence of the solution and provide the regularity of the value function. In Section 6 we establish the convergence of randomized impulse control value function to its classical counterpart as the regularization effect disappears. Based on the randomization framework, we develop an offline RL algorithm in Section 7: we first prove geometric convergence of a policy-improvement scheme, and then introduce a temporal–difference method to implement a model-free version of the RL algorithm, and finally report experimental results on a widely studied linear problem. All the detailed proofs for the theoretical results in Sections 3, 4, 5, 6 and 7 are deferred to Appendix A; additional experimental results are documented in Appendix B.
Notations
We specify the following list of notations for the rest of this paper.
  • Given any $k\in\mathbb{N}$, let $\mathcal{C}^{k}=\mathcal{C}^{k}\left(\mathbb{R}\right)$ be the class of $k$-times-differentiable functions on $\mathbb{R}$. When $k=0$, we use $\mathcal{C}=\mathcal{C}(\mathbb{R})$ to denote the class of continuous functions.
  • Given any $\alpha\in(0,1]$ and any $k\in\mathbb{N}$, let $\mathcal{C}^{k,\alpha}=\mathcal{C}^{k,\alpha}\left(\mathbb{R}\right)$ denote the $\alpha$-Hölder space,
    $\mathcal{C}^{k,\alpha}\coloneqq\left\{f\in\mathcal{C}^{k}:\|f\|_{\mathcal{C}^{k,\alpha}}\coloneqq\|f\|_{\mathcal{C}^{k}}+\left[f^{(k)}\right]_{\alpha}<\infty\right\},$
    where $\|f\|_{\mathcal{C}^{k}\left(K\right)}\coloneqq\sup\limits_{0\leq l\leq k,x\in\mathbb{R}}|f^{(l)}(x)|$ and $\left[f^{(k)}\right]_{\alpha}\coloneq\sup\limits_{x,y\in\mathbb{R},x\neq y}\frac{\left|f^{(k)}(x)-f^{(k)}(y)\right|}{|x-y|^{\alpha}}$, for any function $f\in\mathcal{C}^{k}$ and $f^{(l)}$ being the $l$-th order derivative of $f$ for any $l=0,\dots,k$. Given any precompact subset $K$ of $\mathbb{R}$ denoted by $K\Subset\mathbb{R}$, let $\mathcal{C}^{k,\alpha}\left(K\right)$ denote the $\alpha$-Hölder space restricted to domain $K$. Let $\mathcal{C}^{k,\alpha}_{loc}=\mathcal{C}^{k,\alpha}_{loc}\left(\mathbb{R}\right)$ be the locally $\alpha$-Hölder space,
    $\mathcal{C}^{k,\alpha}_{loc}\coloneqq\left\{f:\mathbb{R}\to\mathbb{R}:f\in\mathcal{C}^{k,\alpha}\left(K\right),\,\forall K\Subset\mathbb{R}\right\}.$
    For simplicity, when $k=0$, we use $\mathcal{C}^{\alpha}$ and $\mathcal{C}^{\alpha}_{loc}$, with $\|\cdot\|_{\mathcal{C}}$ denoted as $\|\cdot\|_{\infty}$.
  • Let $\mathscr{L}$ be the set of Lipschitz-continuous functions,
    $\mathscr{L}\coloneqq\left\{f:\mathbb{R}\to\mathbb{R}:[f]_{1}<\infty\right\}.$
    Denote $\|f\|_{\mathscr{L}}\coloneqq\sup_{x\in\mathbb{R}}\frac{|f(x)|}{1+|x|}$ for any $f\in\mathscr{L}$. Given any $L\in[0,\infty)$, let $\mathscr{L}_{L}$ be the set of $L$-Lipschitz functions,
    $\mathscr{L}_{L}\coloneqq\left\{f\in\mathscr{L}:[f]_{1}\leq L\right\}.$
  • Let $\mathscr{F}$ denote the set of test functions $\phi:\mathbb{R}\to\mathbb{R}$ such that $\inf_{x\in\mathbb{R}}\phi(x)\in\mathbb{R}$ and $\underline{\beta}\coloneqq\inf\{\beta\geq 0:\sup_{x\in\mathbb{R}}\frac{|\phi(x)|}{1+|x|^{\beta}}<+\infty\}<+\infty$.
  • Let $\mathcal{P}\left(\mathbb{R}\right)$ be the set of probability measures on $\mathbb{R}$. For any $p\geq 1$, let $\mathcal{P}_{p}\left(\mathbb{R}\right)$ be the set of probability measures with finite $p$-th moment,
    $\mathcal{P}_{p}\left(\mathbb{R}\right)\coloneqq\left\{\mu\in\mathcal{P}\left(\mathbb{R}\right):\|\mu\|_{p}^{p}\coloneq\int_{\mathbb{R}}|x|^{p}\mu(dx)<\infty\right\}.$
2 Randomized Impulse Control Problem
In this section, we introduce the mathematical formulation of the randomized impulse control problem. Consider a filtered probability space $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0})$ supporting a standard one-dimensional Brownian motion $W=\{W_{t}\}_{t\geq 0}$, where the filtration $\mathbb{F}$ satisfies the usual conditions. If there is no additional specification of probability measures, we will use $\mathbb{E}$ to denote the expectation with respect to $\mathbb{P}$.
On $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$, the uncontrolled state process $X=\{X_{t}\}_{t\geq 0}$ is assumed to satisfy the following dynamic,
$\mathrm{d}X_{t}=b\left(X_{t}\right)\mathrm{d}t+\sigma\left(X_{t}\right)\mathrm{d}W_{t},$ (2.1)
where $X_{0}\sim\mu_{0}\in\mathcal{P}_{2}\left(\mathbb{R}\right)$ such that $X_{0}\perp W$ and $X_{0}\in\mathcal{F}_{0}$. Let $\mathcal{L}$ denote the infinitesimal generator associated with (2.1) such that for any given test function $\phi\in\mathcal{C}^{2}$,
$\mathcal{L}\phi(x)=b(x)\phi^{\prime}(x)+\frac{\sigma(x)^{2}\phi^{\prime\prime}(x)}{2},\quad\forall x\in\mathbb{R}.$ (2.2)
We first specify the following assumptions for the state dynamic (2.1).
Assumption 2.1
The drift $b$ and the volatility $\sigma$ in (2.1) satisfy the following conditions.
  1. (i)
    There exists $L>0$ such that $|b(x)-b(y)|+|\sigma(x)-\sigma(y)|\leq L|x-y|,\quad\forall x,y\in\mathbb{R}.$
  2. (ii)
    There exists $\sigma_{0}>0$ such that $|\sigma(x)|\geq\sigma_{0},\quad\forall x\in\mathbb{R}.$
Under Assumption 2.1, (2.1) admits a strong solution adapted to $\mathbb{F}$ such that $\mathbb{E}[X_{t}^{2}]<\infty$ for all $t\geq 0$; see, for instance, (karatzas1998brownian, , Theorem 2.9). In addition, we have the following technical lemma which is a direct result of applying Itô’s formula, Grönwall’s inequality and Cauchy-Schwarz inequality.
Lemma 2.2
Let $X^{x}=\{X_{t}^{x}\}_{t\geq 0}$ satisfy (2.1) with $X_{0}=x$ for any $x\in\mathbb{R}$. Then,
$\mathbb{E}\left[\sup_{t\geq 0}e^{-\left(L+\frac{1}{2}L^{2}\right)t}|X^{x}_{t}-X^{y}_{t}|\right]\leq|x-y|,\quad\forall x,y\in\mathbb{R}.$
2.1 Review of the Classical Impulse Control Problem
In the literature of the classical impulse control problem without randomization (see, for instance, guo2009smooth ), the objective can be seen as the following optimization problem,
$\psi(x)\coloneqq\inf_{\upsilon\in\Upsilon}J(x;\upsilon),\quad\forall x\in\mathbb{R}.$ (2.3)
Here, $\Upsilon$ denotes the set of admissible impulse control policies, $\upsilon=\bigl\{(\tau_{n},\xi_{n})\bigr\}_{n\geq 1}$, such that $\tau_{n}$ is an $\mathbb{F}$-stopping time and the random jump $\xi_{n}\in\mathcal{F}_{\tau_{n}}$ for any $n\in\mathbb{N}^{+}$, with $\tau_{n}\uparrow\infty$ as $n\to\infty$, $\mathbb{P}$-almost surely. Under any $\upsilon\in\Upsilon$, the controlled state process satisfies
$\mathrm{d}X_{t}=b(X_{t-})\mathrm{d}t+\sigma(X_{t-})\mathrm{d}W_{t}+\sum_{n\geq 1}\xi_{n}\delta(t-\tau_{n}),$ (2.4)
where $X_{0-}\sim\mu_{0}$. Here, $\delta$ denotes the Dirac mass at 0. The cost function $J(x;\upsilon)$ in (2.3) for any $x\in\mathbb{R}$ under any given $\upsilon\in\Upsilon$ is defined as
$J(x;\nu)=\mathbb{E}^{x}\biggl[\int_{0}^{\infty}e^{-rt}f(X_{t-})\mathrm{d}t+\sum_{n\geq 1}e^{-r\tau_{n}}l(\xi_{n})\biggr],$ (2.5)
with discount rate $r>0$, running cost $f:\mathbb{R}\to\mathbb{R}$, and cost of intervention $l:\mathbb{R}\to\mathbb{R}$. We then introduce the following assumptions related to the cost function.
Assumption 2.3
The running cost $f$, the cost of intervention $l$, and the discount rate $r$ satisfy the following conditions.
  1. (i)
    The running cost $f:\mathbb{R}\to[0,+\infty)$ satisfies $f(0)=0$ and there exists $L_{f}>0$ such that $f\in\mathscr{L}_{L_{f}}$, that is,
    $|f(x)-f(y)|\leq L_{f}|x-y|,\quad\forall x,y\in\mathbb{R}.$ (2.6)
  2. (ii)
    The cost of intervention $l:\mathbb{R}\to\mathbb{R}$ satisfies
    1. (a)
      $\exists K>0$ such that $\inf_{x\in\mathbb{R}}l(x)=l(0)=K$ and $l(x)+l(y)\geq l(x+y)+K$ for any $x,y\in\mathbb{R}$;
    2. (b)
      $l(\xi)\to+\infty$ as $|\xi|\to+\infty$;
    3. (c)
      $\exists L_{l}>0$ such that $l\in\mathscr{L}_{L_{l}}$, that is,
      $|l(x)-l(y)|\leq L_{l}|x-y|\quad\forall x,y\in\mathbb{R}.$
  3. (iii)
    Denoting $G\coloneqq L+\frac{1}{2}L^{2}$, the discount factor $r$ satisfies $r-G\in\left(0,\frac{L_{f}}{L_{l}}\right)$.
Let $\mathcal{M}$ be a nonlocal operator such that for any test function $\phi:\mathbb{R}\to\mathbb{R}$,
$\mathcal{M}\phi(x)\coloneqq\inf_{\xi\in\mathbb{R}}\left[\phi(x+\xi)+l(\xi)\right],\quad\forall x\in\mathbb{R}.$ (2.7)
Classical results bensoussan1984impulse ; guo2009smooth show that the value function $\psi$ in (2.3) is the unique viscosity solution of following Hamilton-Jacobi-Bellman (HJB) equation of the quasi-variational-inequality type,
$\min\left\{\mathcal{L}\phi(x)-r\phi(x)+f(x),\mathcal{M}\phi(x)-\phi(x)\right\}=0,\quad\forall x\in\mathbb{R}.$ (HJB-QVI)
Moreover, $\psi\in\mathscr{L}\cap\mathcal{C}^{2}(\mathbb{R}\setminus\partial\mathfrak{C})$, where $\partial\mathfrak{C}$ denotes the boundary of the continuation region $\mathfrak{C}\coloneqq\{x\in\mathbb{R}:\mathcal{L}\psi(x)-r\psi(x)+f(x)=0\}$. The corresponding optimal classical impulse control policy $\upsilon^{*}=\{(\tau^{*}_{n},\xi^{*}_{n})\}_{n\geq 1}$ is given by
$\tau_{1}^{*}=\inf\{t\geq 0:X_{t-}\not\in\mathfrak{C}\};\quad\tau_{n}^{*}=\inf\{t>\tau_{n-1}^{*}:X_{t-}\not\in\mathfrak{C}\},\,n\geq 2.$ (2.8)
$\xi_{n}^{*}=\xi^{*}(X_{\tau_{n}^{*}-})\in\arg\min_{\xi\in\mathbb{R}}\psi(X_{\tau_{n}^{*}-}+\xi)+l(\xi),\quad n\geq 1.$ (2.9)
In (2.9), $\xi^{*}:\mathcal{A}\to\mathbb{R}$ denotes an optimal jump when the current state is in the action region $\mathcal{A}=\mathbb{R}\setminus\mathfrak{C}$.
Apart from characterizing the optimal impulse control policy via (HJB-QVI), an equivalent characterization has also been analyzed in bensoussan1984impulse ; guo2009smooth via the following compound optimal stopping operator $\mathcal{T}$, where for any test function $\phi:\mathbb{R}\to\mathbb{R}$ that is bounded from below,
$\mathcal{T}\phi(x)\coloneqq\inf_{\tau\in\mathbb{T}}\mathbb{E}_{x}\left[\int_{0}^{\tau}e^{-rt}f(X_{t-})\mathrm{d}t+e^{-r\tau}\mathcal{M}\phi(X_{\tau-})\right],$ (classical)
subject to (2.1), where $\mathbb{T}$ denotes the set of $\mathbb{F}$-stopping times such that $\mathbb{P}(\tau<\infty)=1$ for any $\tau\in\mathbb{T}$, and $\mathbb{E}_{x}\left[\cdot\right]=\mathbb{E}\left[\cdot\biggl|X_{0-}=x\right]$. In particular, the value function $\psi$ in (2.3) is a fixed point for $\mathcal{T}$, that is, $\psi=\mathcal{T}\psi$ on $\mathbb{R}$. Correspondingly, from (2.8), we know that $\mathcal{T}\psi(x)$ defined by (classical) admits an optimal stopping time with respect to (2.1),
$\tau^{*}=\inf\{t\geq 0:X_{t-}\not\in\mathfrak{C}\}\in\mathbb{T}.$ (2.10)
2.2 Randomization of Impulse Controls
Let us first re-examine the deterministic nature of the optimal classical impulse control policy in (2.8)–(2.9). and then proceed to the corresponding two-fold randomization scheme for impulse controls. The randomized operators introduced in Sections 2.2.1 and 2.2.2 are the key ingredients for the formulation of the randomized impulse control problem in Section 2.3.
2.2.1 Randomizing the Jump
Deterministic nature of (2.9)
When the current state $x$ is in the action region $\mathcal{A}$, by (HJB-QVI) and the corresponding optimal jump given by (2.9), we must have
$\psi(x)$ $=\mathcal{M}\psi(x)=\inf_{\xi\in\mathbb{R}}\psi(x+\xi)+l(\xi)$
$=\psi(x+\xi^{*}(x))+l(\xi^{*}(x))=\int_{\mathbb{R}}\delta_{\xi^{*}(x)}(\xi)\left[\psi(x+\xi)+l(\xi)\right]\mathrm{d}\xi,$
where $\delta_{y}(\cdot)=\delta(\cdot-y)$ for any $y\in\mathbb{R}$. Let $\mu^{\circ}_{x}:\mathbb{R}\to[0,1]$ be the heaviside function corresponding to $\delta_{\xi^{*}(x)}$ such that $\mu^{\circ}_{x}(y)=\int_{-\infty}^{y}\delta_{\xi^{*}(x)}(\xi)\mathrm{d}\xi,\;\forall y\in\mathbb{R}.$ Then $\mu^{\circ}_{x}$ is the cumulative distribution function for the random variable $Y$ with $\mathbb{P}(Y=\xi^{*}(x))=1$.
Randomization of the jump
For any $x\in\mathbb{R}$, let $\Phi_{x}$ denote the Gaussian distribution $\mathcal{N}(-x,1)$ with mean x-x and variance $1$. Intuitively, $\Phi_{x}$ denotes a simple randomized jump strategy at $x$ such that on average, the jump should send the state back to the origin.
For any $\mu\in\mathcal{P}\left(\mathbb{R}\right)$ such that $\mu\ll\Phi_{x}$, recall the Kullback-Leibler (KL) divergence
$D_{{\rm KL}}\left(\mu\|\Phi_{x}\right)\coloneqq\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho(\zeta)\log\rho(\zeta)\right]\geq 0,$
where $\rho$ is the Radon-Nikodym derivative, $\rho=\frac{d\mu}{d\Phi_{x}}$. Given any $\lambda>0$, define the randomized nonlocal operator $\mathcal{M}^{\lambda}$ such that, for any test function $\phi:\mathbb{R}\to\mathbb{R}$ that is bounded from below,
$\mathcal{M}^{\lambda}\phi(x)\coloneqq\inf_{\mu\in\mathcal{P}\left(\mathbb{R}\right),\mu\ll\Phi_{x}}\int_{\mathbb{R}}\phi(x+\xi)+l(\xi)\mu(\mathrm{d}\xi)+\lambda D_{{\rm KL}}\left(\mu\|\Phi_{x}\right),\quad\forall x\in\mathbb{R}.$ (2.11)
2.2.2 Randomizing the Intervention Time
Deterministic nature of (2.8)
Recall that the value function $\psi$ in (2.3) is a fixed point for the optimal stopping operator defined in (classical). Under $\tau^{*}$ in (2.10), we can rewrite (classical) applied to $\psi$ as follows,
$\mathbb{E}_{x}\left[\int_{0}^{\tau^{*}}e^{-rt}f(X_{t-})\mathrm{d}t+e^{-r\tau^{*}}\mathcal{M}\psi(X_{\tau-})\right]$
$=\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-rt}\left[\mathbbm{1}\left\{\tau^{*}>t\right\}f(X_{t-})+\delta_{\tau^{*}}(t)\mathcal{M}\psi(X_{t-})\right]\mathrm{d}t\right].$
Define the survival process as $p^{\circ}=\{p^{\circ}_{t}\}_{t\geq 0}$ such that $p^{\circ}_{t}=\mathbbm{1}\left\{\tau^{*}>t\right\}$ and $\dot{p}^{\circ}_{t}=\delta_{\tau^{*}}(t)$ for any $t\geq 0$. Intuitively, $p^{\circ}_{t}$ represents the “survival probability” such that the stopping time $\tau^{*}$ is realized after the current time $t$. By construction,
$p^{\circ}_{t}=1-\int_{0}^{t}\dot{p}^{\circ}_{s}\mathrm{d}s,\quad\forall t\geq 0.$
Then we have
$\mathcal{T}\phi(x)=\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-rt}\left(p^{\circ}_{t}f(X_{t-})-\dot{p}^{\circ}_{t}\mathcal{M}\phi(X_{t-})\right)\mathrm{d}t\right].$
Randomization of the intervention time
Similar to dong2024randomized , we borrow the terminology from Cox process jeanblanc2009mathematical with $\pi_{t}\geq 0$ being the intensity for any $t\geq 0$, and define the survival probability by time $t\geq 0$ as $p_{t}=\exp\left\{-\int_{0}^{t}\pi_{s}\mathrm{d}s\right\}$; or equivalently, the process $p=\{p_{t}\}_{t\geq 0}$ satisfies $p_{0}=1$ and
$dp_{t}=\dot{p}_{t}dt=-\pi_{t}p_{t}dt.$ (2.12)
Define $\mathcal{R}:[0,+\infty)\to\mathbb{R}$ as $\mathcal{R}(x)=x-x\log{x}$ for all $x>0$ and $\mathcal{R}(0)=\lim_{x\to 0^{+}}R(x)=0$. Notice that $\mathcal{R}$ attains its maximum at and only at $x=1$. For any $\lambda>0$, define the randomized optimal stopping operator $T^{\lambda}$ such that for any test function $\phi\in\mathscr{F}$ and any $x\in\mathbb{R}$,
$T^{\lambda}\phi(x)=\inf_{\pi=\{\pi_{t}\}_{t\geq 0},\pi_{t}\geq 0}J^{\lambda,\phi}(1,x;\pi)$ (2.13)
subject to (2.1), where for any $p\in[0,1]$,
$J^{\lambda,\phi}(p,x;\pi)\coloneqq\mathbb{E}_{x}\left[\int_{0}^{\infty}pe^{-\int_{0}^{t}r+\pi_{s}\mathrm{d}s}\left(f(X_{t})-\lambda\mathcal{R}(\pi_{t})+\pi_{t}\phi(X_{t})\right)\mathrm{d}t\right],$ (2.14)
for any $\pi=\{\pi_{t}\}_{t\geq 0}$ with $\pi_{t}\geq 0$ for all $t\geq 0$. For simplicity, we abuse the notation slightly and denote $J^{\lambda,\phi}(\cdot;\pi)=J^{\lambda,\phi}(1,\cdot;\pi)$ for any intensity process $\pi$.
2.3 The Two-Fold Randomization of Impulse Controls
Analogous to (classical), for the randomized impulse control problem, we define the following compound optimal stopping operator $\mathcal{T}^{\bm{\lambda}}$ under any given $\bm{\lambda}=(\lambda_{1},\lambda_{2})\in(0,+\infty)^{2}$ such that for any test function $\phi\in\mathscr{F}$ and any $x\in\mathbb{R}$ and subject to (2.1),
$\mathcal{T}^{\bm{\lambda}}\phi(x)=$ $T^{\lambda_{1}}\left[\mathcal{M}^{\lambda_{2}}\phi\right](x)$ (randomized)
$=$ $\inf_{\pi}\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}r+\pi_{s}\mathrm{d}s}\left(f(X_{t})-\lambda_{1}\mathcal{R}(\pi_{t})+\pi_{t}\mathcal{M}^{\lambda_{2}}\phi(X_{t})\right)\mathrm{d}t\right].$
In light of the fixed-point characterization of the classical impulse control problem via (classical), we have the following definition of the solution to the randomized impulse control problem via (randomized).
Definition 2.4
A function $\psi^{\bm{\lambda}}:\mathbb{R}\to\mathbb{R}$ is said to be a value function for the randomized impulse control problem associated with $\mathcal{T}^{\bm{\lambda}}$ in (randomized) under any given $\bm{\lambda}=(\lambda_{1},\lambda_{2})\in(0,+\infty)^{2}$, if $\psi^{\bm{\lambda}}\in\mathscr{L}$ is bounded from below, and $\psi^{\bm{\lambda}}=\mathcal{T}^{\bm{\lambda}}\psi^{\bm{\lambda}}$ on $\mathbb{R}$. Moreover, if $\pi^{*}=\{\pi^{*}_{t}\}_{t\geq 0}$, with $\pi^{*}_{t}\geq 0$ for all $t\geq 0$, and $\mu^{*}=\{\mu^{*}_{x}\}_{x\in\mathbb{R}}$, with $\mu^{*}_{x}\in\{\mu\in\mathcal{P}\left(\mathbb{R}\right):\mu\ll\Phi_{x}\}$ for all $x\in\mathbb{R}$, satisfy
$\begin{cases}\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)=\mathbb{E}_{\xi\sim\mu^{*}_{x}}\left[\psi^{\bm{\lambda}}(x+\xi)+l(\xi)\right]+\lambda_{2}D_{{\rm KL}}\left(\mu^{*}_{x}\|\Phi(x)\right),\\ \psi^{\bm{\lambda}}(x)=\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}r+\pi^{*}_{s}\mathrm{d}s}\left(f(X_{t})-\lambda_{1}\mathcal{R}(\pi^{*}_{t})+\pi^{*}_{t}\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(X_{t})\right)\mathrm{d}t\right],\end{cases}\;\;\forall x\in\mathbb{R},$
then we call $\upsilon^{\bm{\lambda}}=(\pi^{*},\mu^{*})$ an optimal randomized impulse control policy associated with $\psi^{\bm{\lambda}}$.
3 Preliminary Analysis on the Randomized Operators
In this section, we first study the properties of the randomized operators introduced in Section 2. These properties are fundamental for developing the HJB characterization for the randomized problem in Section 4, and for solving the randomized impulse control problem according to Definition 2.4 in Section 5.
3.1 The Randomized Nonlocal Operator
First, we introduce several basic properties of the randomized nonlocal operator $\mathcal{M}^{\lambda}$ in (2.11) being applied to test functions $\phi\in\mathscr{F}$.
Lemma 3.1
Fix any $\lambda>0$.
  1. (i)
    For any $\phi\in\mathscr{F}$, $\mathcal{M}^{\lambda}\phi(x)\in\mathbb{R}$ for all $x\in\mathbb{R}$.
  2. (ii)
    For any $\phi_{1},\phi_{2}\in\mathscr{F}$, $\left\|\mathcal{M}^{\lambda}\phi_{1}-\mathcal{M}^{\lambda}\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}\leq\left\|\phi_{1}-\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}$.
  3. (iii)
    $\mathcal{M}^{\lambda}$ is concave, i.e., for any $\phi_{1},\phi_{2}\in\mathscr{F}$ and any $\eta\in[0,1]$,
    $\mathcal{M}^{\lambda}\left[\eta\phi_{1}+(1-\eta)\phi_{2}\right]\geq\eta\mathcal{M}^{\lambda}\phi_{1}+(1-\eta)\mathcal{M}^{\lambda}\phi_{2}.$
  4. (iv)
    $\mathcal{M}^{\lambda}$ is nondecreasing, i.e., for any $\phi_{1},\phi_{2}\in\mathscr{F}$ with $\phi_{1}\leq\phi_{2}$ on $\mathbb{R}$, $\mathcal{M}^{\lambda}\phi_{1}\leq\mathcal{M}^{\lambda}\phi_{2}$.
The next results concern with $\mathcal{M}^{\lambda}$ being applied to more restricted families of test functions, in particular, Lipschitz test functions. This additional restriction is also in alignment with the Lipschitz value function introduced in Definition 2.4.
Lemma 3.2
Fix any $\lambda>0$ and any test function $\phi,\phi_{1},\phi_{2}\in\mathscr{L}$ that is bounded from below. Then, $\mathcal{M}^{\lambda}\phi\in\mathscr{L}_{L_{l}}$, and
$\mathcal{M}^{\lambda}\phi(x)=-\lambda\log\mathbb{E}_{\zeta\sim\Phi_{x}}\left[e^{-\frac{\phi(x+\zeta)+l(\zeta)}{\lambda}}\right],\quad\forall x\in\mathbb{R}.$ (3.1)
Moreover, there exist $\kappa_{1},\kappa_{2}>0$ with which we define $\delta_{t}\coloneqq 2\kappa_{1}e^{-\kappa_{2}t^{2}}$ for all $t\geq 0$ and
${\rm subG}_{x}\left(\kappa_{1},\kappa_{2}\right)\coloneqq\left\{\nu\in\mathcal{P}\left(\mathbb{R}\right):\mathbb{P}(|x+\xi|>t)\leq\delta_{t},\,\xi\sim\nu,\,\forall t\geq 0\right\},$ (3.2)
such that
$\mathcal{M}^{\lambda}\phi(x)=\mathbb{E}_{\xi\sim\mu_{x}}\left[\phi(x+\xi)+l(\xi)\right]+\lambda D_{{\rm KL}}\left(\mu\|\Phi_{x}\right),\quad\forall x\in\mathbb{R},$
where $\mu_{x}\in{\rm subG}_{x}\left(\kappa_{1},\kappa_{2}\right)$ with $\frac{d\mu_{x}(\xi)}{d\Phi_{x}(\xi)}\propto e^{-\frac{\phi(x+\xi)+l(\xi)}{\lambda}}$.
Furthermore, for any $R>0$, there exists $\mathscr{E}_{R}>0$ with $\lim_{R\to\infty}\mathscr{E}_{R}=0$ such that
$\|\mathcal{M}^{\lambda}\phi_{1}-\mathcal{M}^{\lambda}\phi_{2}\|_{L^{\infty}(\mathbb{R})}\leq\|\phi_{1}-\phi_{2}\|_{L^{\infty}([-R,R])}+\mathscr{E}_{R}.$
Finally, we have the following comparison results among $\mathcal{M}^{\lambda}$’s with various $\lambda>0$ and $\mathcal{M}$.
Lemma 3.3
The properties hold as follows:
  1. (i)
    For any test function $\phi\in\mathscr{L}$ that is bounded from below, $\mathcal{M}^{\lambda}\phi\leq\mathcal{M}^{\lambda^{\prime}}\phi$ for all $\lambda,\lambda^{\prime}\in(0,+\infty)$ such that $\lambda<\lambda^{\prime}$.
  2. (ii)
    For any $\lambda>0$ and any $\phi\in\mathscr{F}$, $\mathcal{M}\phi\leq\mathcal{M}^{\lambda}\phi$ on $\mathbb{R}$.
  3. (iii)
    For any test function $\phi\in\mathscr{L}$ that is bounded from below, $\lim_{\lambda\to 0^{+}}\|\mathcal{M}^{\lambda}\phi-\mathcal{M}\phi\|_{L^{\infty}(K)}=0$ for any $K\Subset\mathbb{R}$.
3.2 The Randomized Optimal Stopping Operator
In this section, we focus on the randomized optimal stopping operator $T^{\lambda}$ being applied to any Lipschitz test function $\phi$. Through a corresponding HJB characterization and a verification theorem, we show that $T^{\lambda}\phi$ is the value function attained at a unique optimal intensity process $\pi^{*}$.
Lemma 3.4
For any $\lambda>0$, any $L_{\phi}>0$, and any $\phi\in\mathscr{L}_{L_{\phi}}$ that is bounded from below, $T^{\lambda}\phi\in\mathscr{L}_{\max\{\frac{L_{f}}{r-G},L_{\phi}\}}$.
Note that when $\phi\in\mathscr{L}$ bounded from below, the following conventional optimal control problem is well-posed, where for any $x\in\mathbb{R}$ and any $p\in[0,1]$,
$\left\{\begin{aligned} v^{\lambda,\phi}(p,x)=&\inf_{\pi}\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}p_{t}\left(f(X_{t})-\lambda\mathcal{R}(\pi_{t})+\pi_{t}\phi(X_{t})\right)\mathrm{d}t\right]\\ \text{subject to }\mathrm{d}X_{t}=&b(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}W_{t},\ \ X_{0}=x,\\ \mathrm{d}p_{t}=&-\pi_{t}p_{t}\mathrm{d}t,\ \ p_{0}=p.\end{aligned}\right$. (3.3)
Moreover, we observe that for any $x\in\mathbb{R}$, $T^{\lambda}\phi(x)=v^{\lambda,\phi}(1,x)$. By the dynamic programming principle, the HJB associated with (3.3) is given by
$(\mathcal{L}-r)v(p,x)+pf(x)+p\inf_{\pi\geq 0}\left\{\pi(x)\left[\phi(x)-\partial_{p}v(p,x)\right]-\lambda\mathcal{R}(\pi(x))\right\}=0,$ (3.4)
for $v:[0,1]\times\mathbb{R}\to\mathbb{R}$, $\forall(p,x)\in(0,1)\times\mathbb{R}$. Observe that $v^{\lambda,\phi}(p,x)=pT^{\lambda}\phi(x)$. Then from (3.4), we deduce the HJB associated with $T^{\lambda}\phi(x)$ as follows,
$0=(\mathcal{L}-r)v+f+\inf_{\pi\geq 0}\left\{\pi(\phi-v)-\lambda\mathcal{R}(\pi)\right\}\\ =(\mathcal{L}-r)v+f-\lambda e^{-\frac{\phi-v}{\lambda}},$ (3.5)
for $v:\mathbb{R}\to\mathbb{R}$, and the infimum is attained at $\bar{\pi}(x)=e^{-\frac{\phi(x)-v(x)}{\lambda}}$ for any $x\in\mathbb{R}$, due to the strong concavity of $\mathcal{R}$.
We have the following verification theorem, establishing $T^{\lambda}\phi$ as a solution to the above HJB.
Theorem 3.5
Fix any $\lambda>0$ and any $\phi\in\mathscr{L}$ that is bounded from below, the value function $T^{\lambda}\phi$ in (2.13) is the unique classical solution to the HJB equation (3.5) in $\mathcal{C}^{2}\cap\mathscr{L}$. Moreover, the intensity process $\pi^{*}=\{\pi^{*}_{t}\}_{t\geq 0}$ with $\pi^{*}_{t}=\bar{\pi}(X_{t})$ for all $t\geq 0$, where
$\bar{\pi}(x)=e^{-\frac{\phi(x)-T^{\lambda}\phi(x)}{\lambda}},\qquad\forall x\in\mathbb{R},$ (3.6)
is the corresponding optimal control policy such that
$T^{\lambda}\phi(x)=J^{\lambda,\phi}(x;\pi^{*}),\quad\forall x\in\mathbb{R}.$
3.3 The Compound Optimal Stopping Operator
Composed with $\mathcal{M}^{\lambda_{2}}$ and $T^{\lambda_{1}}$, the properties presented in Sections 3.1 and 3.2 lead to the following properties of $\mathcal{T}^{\bm{\lambda}}$. These properties lay the foundation for the validity of Definition 2.4.
Lemma 3.6
The compound optimal stopping operator $\mathcal{T}^{\bm{\lambda}}$ satisfies the following properties:
  1. (i)
    $\mathcal{T}^{\bm{\lambda}}$ is non-decreasing, i.e., for any $\phi_{1},\phi_{2}\in\mathscr{F}$ with $\phi_{1}\leq\phi_{2}$, $\mathcal{T}^{\bm{\lambda}}\phi_{1}\leq\mathcal{T}^{\bm{\lambda}}\phi_{2}$.
  2. (ii)
    For any $\bm{\lambda}\in(0,+\infty)^{2}$ and any test function $\phi\in\mathscr{L}$ that is bounded from below, $\mathcal{T}^{\bm{\lambda}}\phi\in\mathscr{L}_{\frac{L_{f}}{r-G}}$.
  3. (iii)
    For any $\bm{\lambda}\in(0,+\infty)^{2}$ and any test function $\phi_{1},\phi_{2}\in\mathscr{L}$ that are bounded from below, $\|\mathcal{T}^{\bm{\lambda}}\phi_{1}-\mathcal{T}^{\bm{\lambda}}\phi_{2}\|_{L^{\infty}(\mathbb{R})}\leq\|\phi_{1}-\phi_{2}\|_{L^{\infty}(\mathbb{R})}$.
Lemma 3.6-(i) directly follows from Lemma 3.1-(iv) and its proof is therefore omitted.
4 HJB Characterization of Randomized Impulse Control Problem and Verification Theorem
In this section, we provide the HJB characterization (HJB) for the randomized impulse control solution in Definition 2.4 under a fixed randomization parameter $\bm{\lambda}=(\lambda_{1},\lambda_{2})\in(0,+\infty)^{2}$. With the help of an alternative randomization scheme (4.9), we then provide a verification theorem for (HJB), which leads to the uniqueness of solution in Definition 2.4.
4.1 HJB Characterization
According to the compound optimal stopping operator defined in (randomized), we have $\mathcal{T}^{\bm{\lambda}}\phi=T^{\lambda_{1}}\left[\mathcal{M}^{\lambda_{2}}\phi\right]$ for any $\phi\in\mathscr{F}$. Following (3.5) and (3.6), we can derive the HJB equation associated with $T^{\lambda_{1}}\left[\mathcal{M}^{\lambda_{2}}\phi\right]$,
$0=(\mathcal{L}-r)v+f+\inf_{\pi\geq 0}\left\{\pi\left(\mathcal{M}^{\lambda_{2}}\phi-v\right)-\lambda_{1}\mathcal{R}(\pi)\right\}\\ =(\mathcal{L}-r)v+f-\lambda_{1}e^{-\frac{\mathcal{M}^{\lambda_{2}}\phi-v}{\lambda_{1}}},$ (4.1)
for function $v:\mathbb{R}\to\mathbb{R}$. Accordingly, the optimal feedback policy $\pi^{*}$ is given by
$\pi^{*}(x)=e^{-\frac{\mathcal{M}^{\lambda_{2}}\phi-\mathcal{T}^{\bm{\lambda}}\phi}{\lambda_{1}}}.$ (4.2)
By Definition 2.4, if a function $\psi^{\bm{\lambda}}:\mathbb{R}\to\mathbb{R}$ is a value function for the randomized impulse control problem, then it must satisfy the fixed-point condition,
$\psi^{\bm{\lambda}}=\mathcal{T}^{\bm{\lambda}}\psi^{\bm{\lambda}}=T^{\lambda_{1}}\left[\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}\right].$
Thus, by (4.1), we have the following HJB equation associated with $\psi^{\bm{\lambda}}$,
$(\mathcal{L}-r)v+f-\lambda_{1}e^{-\frac{\mathcal{M}^{\lambda_{2}}v-v}{\lambda_{1}}}=0,$ (HJB)
for $v:\mathbb{R}\to\mathbb{R}$. From (4.2), we know that the optimal feedback policy associated with $\psi^{\bm{\lambda}}$ is
$\pi^{*}(x)=e^{-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}-\psi^{\bm{\lambda}}}{\lambda_{1}}}.$ (4.3)
Theorem 4.1
Suppose a function $V:\mathbb{R}\to\mathbb{R}$ is a value function for the randomized impulse control problem according to Definition 2.4, then $V\in\mathcal{C}^{2}$ is a classical solution to (HJB).
4.2 Verification Theorem
We will provide a verification theorem for (HJB) through an alternative randomization scheme of impulse control.
4.2.1 Connection with Randomization via Controlled Compound Poisson Process
In this section, we discuss an alternative approach to introduce randomness to the classical impulse control policy under some fixed $\bm{\lambda}=(\lambda_{1},\lambda_{2})\in(0,+\infty)$.
Recall the optimal impulse control policy characterized by (2.8)–(2.9). We can write the corresponding control process as a jump process $Y^{*}=\{Y^{*}_{t}\}_{t\geq 0}$, where
$Y^{*}_{t}=\sum_{n=1}^{N^{*}_{t}}\xi^{*}_{n},\ \ \text{where}\ \ N^{*}_{t}=\sum_{n\geq 1}^{\infty}\mathbbm{1}\{\tau_{n}^{*}\leq t\},\quad\forall t\geq 0.$ (4.4)
Here, $\{(\tau^{*}_{n},\xi^{*}_{n})\}_{n\geq 1}$ is given by (2.8)–(2.9). The deterministic nature of $Y^{*}$ is characterized as follows: in (4.4), the associated counting process $N^{*}$ is represented by the sequence of arrival times $\{\tau^{*}_{n}\}_{n\geq 1}$, which is of a deterministic feedback form; the sequence of jumps $\{\xi^{*}_{n}\}_{n\geq 1}$ is also of a deterministic feedback form.
Inspired by denkert2025control , we embed the impulse mechanism via a controlled compound Poisson process. Consider $\pi=\{\pi_{t}\}_{t\geq 0}$ such that $\pi_{t}\geq 0$ for all $t\geq 0$, and let $N^{\pi}=\{N^{\pi}_{t}\}_{t\geq 0}$ be a nonhomogeneous Poisson process with intensity $\pi$ such that, conditioning on $\pi$, $N^{\pi}\perp W$. Let $\eta=\{\eta_{n}\}_{n=1}^{\infty}$ be a sequence of independent jumps such that $\eta\perp(W,N^{\pi})$. Define the following jump process $Y^{\pi,\bm{\mu}}=\{Y^{\pi,\bm{\mu}}_{t}\}_{t\geq 1}$ such that
$Y^{\pi,\bm{\mu}}_{t}=\sum_{n=1}^{N^{\pi}_{t}}\eta_{n},\qquad\forall t\geq 0.$ (4.5)
Here, $\bm{\mu}=\{\mu_{t}\}_{t\geq 0}$ is a flow of probability measures in $\mathcal{P}\left(\mathbb{R}\right)$ such that for any open set $U\subset\mathbb{R}$,
$\mathbb{P}(Y^{\pi,\bm{\mu}}_{t}\in U)=\int_{0}^{t}Y^{\pi,\bm{\mu}}(\mathrm{d}s,\mathrm{d}\xi)=\int_{0}^{t}\mu_{s}(\mathrm{d}\xi)\pi_{s}\mathrm{d}s,\quad\forall t\geq 0.$
We then consider the following controlled state process,
$X_{t}=X_{0-}+\int_{0}^{t}b(X_{s-})\mathrm{d}s+\int_{0}^{t}\sigma(X_{s-})\mathrm{d}W_{s}+\int_{0}^{t}\int_{\mathbb{R}}\xi Y^{\pi,\bm{\mu}}(\mathrm{d}s,\mathrm{d}\xi),\quad\forall t\geq 0.$ (4.6)
The infinitesimal generator with respect to (4.6) is given by
$\hat{\mathcal{L}}^{\pi,\bm{\mu}}\phi(t,x)$ $=\partial_{t}\phi(t,x)+\mathcal{L}\phi(t,x)+\pi_{t}\mathbb{E}_{\xi\sim\mu_{t}}\left[\phi(t,x+\xi)-\phi(t,x)\right],$ (4.7)
for any test function $\phi:[0,+\infty)\times\mathbb{R}$, $(t,x)\in(0,+\infty)\times\mathbb{R}$.
Let $\mathcal{A}$ denote the set of admissible pairs $(\pi,\bm{\mu})$ such that for all $t\geq 0$, $\pi_{t}=\pi(t,X_{t-})$ for some $\pi:[0,+\infty)\times\mathbb{R}\to[0,+\infty$, and $\mu_{t}=\bm{\mu}[t,X_{t-}]$ such that $\bm{\mu}[t,y]\in\mathcal{P}\left(\mathbb{R}\right)$ with $\bm{\mu}[t,y]\ll\Phi_{y}$ for all $(t,y)\in[0,+\infty)\times\mathbb{R}$; for such $\bm{\mu}$, let $\bm{\varrho}[t,y]=\frac{d\bm{\mu}[t,y]}{d\Phi_{y}}$ and denote $\bm{\varrho}=\{\rho_{t}\}_{t\geq 0}$ with $\rho_{t}=\bm{\varrho}[t,X_{t-}]$, for all $t\geq 0$ and $y\in\mathbb{R}$. Later we may use $\bm{\mu}$ and $\bm{\varrho}$ interchangeably. For any $(\pi,\bm{\mu})\in\mathcal{A}$, define the cost function $J^{\bm{\lambda}}$ with respect to the state process (4.6) as follows,
$J^{\bm{\lambda}}(x;\pi,\bm{\mu})$ $=\mathbb{E}_{x}\Big[\int_{0}^{\infty}e^{-rt}\Big\{f(X_{t})-\lambda_{1}\mathcal{R}(\pi_{t})$
$\hskip 50.0pt+\int_{\mathbb{R}}\left[l(\xi)-\lambda_{2}\mathcal{H}(\mu_{t};X_{t-})\right]Y^{\pi,\bm{\mu}}(\mathrm{d}t,\mathrm{d}\xi)\Big\}\mathrm{d}t\Big],\quad\forall x\in\mathbb{R},$
with $\mathcal{H}(\nu;x)=-D_{KL}(\nu\|\Phi_{x})$ for any $\nu\in\mathcal{P}\left(\mathbb{R}\right)$ such that $\nu\ll\Phi_{x}$. Observe that $\mathcal{R}$ and $\mathcal{H}$ serve as the entropy regularization to encourage randomization of the counting process and sequence of jumps, respectively. We are interested in the optimal control problem,
$\tilde{\psi}^{\bm{\lambda}}(x)=\inf_{(\pi,\bm{\mu})\in\mathcal{A}}J^{\bm{\lambda}}(x;\pi,\bm{\mu}),$ (4.8)
subject to (4.6) with $X_{0-}=x$ for any $\forall x\in\mathbb{R}$.
Notice that $D_{{\rm KL}}\left(\bm{\mu}[t,y]\|\Phi_{y}\right)=\mathbb{E}_{\zeta\sim\Phi_{y}}\left[\bm{\varrho}[t,y](\zeta)\log{\bm{\varrho}[t,y](\zeta)}\right]$ for all $t\geq 0$ and $y\in\mathbb{R}$. Then, the cost function can be rewritten as
$J^{\bm{\lambda}}(x;\pi,\bm{\mu})=$ $\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-rt}\left\{f(X_{t})-\lambda_{1}\mathcal{R}(\pi_{t})+\pi_{t}\left[\mathbb{E}_{\xi\sim\mu_{t}}\left[l(\xi)\right]-\lambda_{2}\mathcal{H}(\mu_{t};X_{t-})\right]\right\}\mathrm{d}t\right]$
$=$ $\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-rt}\left(f(X_{t})-\lambda_{1}\mathcal{R}(\pi_{t})+\pi_{t}\mathbb{E}_{\zeta\sim\Phi_{X_{t-}}}\left[\rho_{t}(\zeta)\left(l(\zeta)+\lambda_{2}\log{\rho_{t}(\zeta)}\right)\right]\right)\mathrm{d}t\right].$
Therefore, the optimal control problem (4.8) is equivalent to
$\tilde{\psi}^{\bm{\lambda}}(x)$ $=\inf_{(\pi,\rho)\in\mathcal{A}}\mathbb{E}\Big[\int_{0}^{\infty}e^{-rt}\Big(f(X_{t})-\lambda_{1}\mathcal{R}(\pi_{t})+\pi_{t}\mathbb{E}_{\xi\sim\Phi_{X_{t-}}}\left[\rho_{t}(\xi)\left(l(\xi)+\lambda_{2}\log{\rho_{t}(\xi)}\right)\right]\Big)\mathrm{d}t\Big],$ (4.9)
subject to (4.6) with $X_{0-}=x$ for any $x\in\mathbb{R}$.
We have the following equivalence between the fixed-point characterization of randomized impulse control given by Definition (2.4) and the control problem (4.9) via their associated HJB’s.
Lemma 4.2
The HJB equation associated with (4.9) coincides with (HJB).
4.2.2 Verification Theorem
With the alternative randomization framework (4.9) and the equivalence relation given by Lemma 4.2, we present the following verification theorem for (HJB).
Theorem 4.3
Suppose $V\in\mathcal{C}^{2}\cap\mathscr{L}$ is a classical solution to (HJB) that is bounded from below. Then $V=\tilde{\psi}^{\bm{\lambda}}$ as in (4.9).
Theorem 4.1 and 4.3 immediately lead to the following uniqueness result.
Theorem 4.4
There exists at most one value function $\psi^{\bm{\lambda}}\in\mathscr{L}$ with a finite lower bound according to Definition 2.4.
Theorem 4.4 guarantees that the solution to the randomized impulse control framework introduced in Definition 2.4 is free from ambiguity.
5 Solving Randomized Impulse Control Problem via an Iterative Approach
In this section, we provide a constructive proof for the existence of solution to the randomized impulse control problem specified in Definition 2.4 under a fixed $\bm{\lambda}\in(0,+\infty)^{2}$. Its building block is an iterative approach mirroring the classical Howard iteration howard1960dynamic . The construction of the iterates consists of two key components, namely the “policy evaluation” and the “policy improvement”. Tailored to our randomized impulse control problem, these two components are achieved through the compound optimal stopping operator $\mathcal{T}^{\bm{\lambda}}$ defined in (randomized). We show that this iterative approach leads to a limiting function, which is indeed a value function under Definition 2.4.
5.1 Construction of the iterates
At the initial step $n=0$, we adopt a conservative policy, simply not intervening at all. That is, $\pi^{0}_{t}\equiv 0$ for all $t\geq 0$. Correspondingly, the total cost $\psi^{\bm{\lambda},0}:\mathbb{R}\to\mathbb{R}$ is defined as
$\psi^{\bm{\lambda},0}(x)=\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(X_{t})\mathrm{d}t\Big|X_{0}=x\right],\quad\forall x\in\mathbb{R}.$ (5.1)
Starting with $\psi^{\bm{\lambda},0}$, we continue the construction recursively via $\mathcal{T}^{\bm{\lambda}}$, namely for any $n\in\mathbb{N}^{+}$,
$\psi^{\bm{\lambda},n}(x)$ $=\mathcal{T}^{\bm{\lambda}}\psi^{\bm{\lambda},n-1}(x)$ (5.2)
$=\inf_{\pi}\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-rt}p_{t}\left(f(X_{t})+\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(X_{t})\pi_{t}-\lambda_{1}\mathcal{R}(\pi_{t})\right)\mathrm{d}t\right],$
subject to (2.1) and (2.12) with $p_{0}=1$. Then we have the following property for the iterates $\psi^{\bm{\lambda},n}$’s.
Lemma 5.1
For any $n\in\mathbb{N}$, $\psi^{\bm{\lambda},n}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$.
From Lemma 5.1, we know that $\psi^{\bm{\lambda},n},\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$ with a natural lower bound 0 for all $n\in\mathbb{N}$. Therefore by Theorem 3.5, for $n\geq 1$, $\psi^{\bm{\lambda},n}$ in (5.2) is the unique classical solution in $\mathcal{C}^{2}$ to the following HJB equation,
$0$ $=\mathcal{L}\psi^{\bm{\lambda},n}-r\psi^{\bm{\lambda},n}+f+\inf_{\pi}\left\{\left(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}-\psi^{\bm{\lambda},n}\right)\pi-\lambda_{1}\mathcal{R}(\pi)\right\}$ (5.3)
$=\mathcal{L}\psi^{\bm{\lambda},n}-r\psi^{\bm{\lambda},n}+f-\lambda_{1}e^{-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)}{\lambda_{1}}}.$
In addition, the corresponding optimal policy $\pi^{n}$ is given by
$\pi^{n}(x)=e^{-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)}{\lambda_{1}}};$ (5.4)
denote $p^{n}_{t}=e^{-\int_{0}^{t}\pi^{n}_{s}\mathrm{d}s}$ for any $t\geq 0$.
5.2 Convergence of the Iterates
The next result shows that adopting the policy in (5.4) is indeed a policy improvement step, and correspondingly, the policy evaluation step according to (5.2) will eventually converge to a Lipschitz-continuous function.
Theorem 5.2
For $\{\psi^{\bm{\lambda},n}\}_{n\geq 0}$ constructed in Section 5.1,
  1. (i)
    $\psi^{\bm{\lambda},n+1}\leq\psi^{\bm{\lambda},n}$ for all $n\in\mathbb{N}$;
  2. (ii)
    $\exists\hat{\psi}^{\bm{\lambda}}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$ such that $\psi^{\bm{\lambda},n}\xrightarrow{n\to\infty}\hat{\psi}^{\bm{\lambda}}$ pointwise. Moreover, $\psi^{\bm{\lambda},n}$ converges to $\hat{\psi}^{\bm{\lambda}}$ locally uniformly.
Next, we show that such a pointwise limit $\hat{\psi}^{\bm{\lambda}}$ actually corresponds to the randomized impulse control value function.
Theorem 5.3
The pointwise limit $\hat{\psi}^{\bm{\lambda}}$ in Theorem 5.2 is a value function for the randomized impulse control problem as in Definition 2.4. The associated optimal policy $(\pi^{*}=\{\pi^{*}_{t}\}_{t\geq 0},\mu^{*}=\{\mu^{*}_{x}\}_{x\in\mathbb{R}})$ is given by $\mu^{*}_{x}\in\mathcal{P}\left(\mathbb{R}\right)$ such that
$\rho^{*}_{x}(\xi)=\frac{d\mu^{*}_{x}(\xi)}{d\Phi_{x}(\xi)}=\frac{e^{-\frac{\hat{\psi}^{\bm{\lambda}}(x+\xi)+l(\xi)}{\lambda_{2}}}}{\mathbb{E}_{\zeta\sim\Phi_{x}}\left[e^{-\frac{\hat{\psi}^{\bm{\lambda}}(x+\zeta)+l(\zeta)}{\lambda_{2}}}\right]},\quad\forall x\in\mathbb{R},$
and $\pi^{*}_{t}=\bar{\pi}(X_{t})$ for any $t\geq 0$ with $\bar{\pi}=e^{-\frac{\mathcal{M}^{\lambda_{2}}\hat{\psi}^{\bm{\lambda}}-\hat{\psi}^{\bm{\lambda}}}{\lambda_{1}}}.$
Theorem 5.2 and 5.3 guarantee that the iterative approach with initialization (5.1) and recursive relation (5.2) indeed converges to the value function of the randomized impulse control problem according to Definition 2.4. Together with the uniqueness result in Theorem 4.4, we conclude that the randomized impulse control problem via the compound optimal stopping operator (randomized) is well-defined for any given $\bm{\lambda}\in(0,+\infty)^{2}$.
6 Convergence of the Randomized Problem to the Classical Problem
In Sections 4 and 5, we have shown the well-definedness of the randomized impulse control problem in Definition 2.4 for any given randomization parameter $\bm{\lambda}$. In this section, we focus on the relation between this randomized problem and the classical impulse control problem (2.3). In particular, we establish a convergence result for the randomized impulse–control problem to its classical counterpart as $\bm{\lambda}$ approaches $(0,0)$.
Let us first introduce the sequence of functions $\{\psi^{n}\}_{n\geq 0}$, representing the classical policy iterates. This sequence converges to the classical solution $\psi$, which solves (HJB-QVI). Assume that $\psi^{0}$ is the value function where no intervention is made. Namely, $\psi^{0}\equiv\psi^{\bm{\lambda},0}$ as defined in (5.1). Starting with $\psi^{0}$, we recursively construct the iteration with $\mathcal{T}$ (recall (classical)). For any $n\in\mathbb{N}^{+}$,
$\psi^{n}(x)=\mathcal{T}\psi^{n-1}(x)=\inf_{\tau}\mathbb{E}_{x}\left[\int_{0}^{\tau}e^{-rt}f(X_{t})\mathrm{d}t+\mathcal{M}\psi^{n-1}(X_{t})\right],$ (6.1)
subject to (2.1).
Equivalently, one can show by classical results bensoussan2011applications ; pham2009continuous that when $n\in\mathbb{N}^{+}$, $\psi^{n}$’s are the unique viscosity solution to the corresponding HJB equation:
$\min\{\mathcal{L}\psi^{n}-r\psi^{n}+f,\mathcal{M}\psi^{n-1}-\psi^{n}\}=0.$ (6.2)
We first state the following assumption, which will be active from now on.
Assumption 6.1
We impose the following boundedness conditions:
  1. (i)
    The coefficients satisfy $\|b\|_{L^{\infty}(\mathbb{R})}+\|\sigma\|_{L^{\infty}(\mathbb{R})}+\|f\|_{L^{\infty}(\mathbb{R})}\leq M$.
  2. (ii)
    The continuation regions $\mathfrak{C}^{n}$ of the policy iterates $\{\psi^{n}\}_{n\geq 1}$ are uniformly bounded; that is, there exists a bounded set $W\Subset\mathbb{R}$ such that $\mathfrak{C}^{n}\subset W$ for all $n$.
The first condition is a common convention in the literature on impulse control problems reisinger2019penalty ; reisinger2020error . The second condition is satisfied in many practical settings. For instance, both our example in Section 7.2 and the numerical study in reisinger2020error meet this requirement.
Under Assumption 6.1, one can show that the classical policy iterates $\{\psi^{n}\}_{n\geq 0}$ locally converges to the classical solution $\psi$ at a geometric rate.
Proposition 6.2 ((reisinger2019penalty, , Theorem 3.4) and (reisinger2020error, , Proposition 4.1))
For any $n\in\mathbb{N}^{+}$, let $\psi$ and $\psi^{n}$ be the unique viscosity solution to (HJB-QVI) and (6.2) respectively. Then there exist constants $\mu\in(0,1]$ and $C>0$ such that
$0\leq\psi^{n}-\psi\leq C(1-\mu)^{n},\quad n\in\mathbb{N}.$
Consequently, the iterates $\{\psi^{n}\}_{n\geq 0}$ are bounded uniformly in $n$.
Given the similarity in the problem settings, the proof of Proposition 6.2 is essentially the same as that of (reisinger2020error, , Proposition 4.1), therefore omitted.
Also, one can prove that the iterates $\{\psi^{n}\}_{n\geq 0}$ have a uniform Lipschitz constant.
Lemma 6.3
For any $n\in\mathbb{N}$, $\psi^{n}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$.
To better analyze the fixed point iteration of the classical stopping operator, we introduce a “semi-randomized” impulse control problem, where we merely randomize the jump size with $\mathcal{M}^{\lambda_{2}}\psi^{n}$. In specific, for any $n\in\mathbb{N}$,
$\begin{cases}\mathcal{M}^{\lambda_{2}}\phi(x)&=\inf_{\mu\in\mathcal{P}(\mathbb{R})}\left[\int_{\mathbb{R}}\phi(x+\xi)+l(\xi)\mu(d\xi)-\lambda_{2}\mathcal{H}(\mu)\right],\\ \tilde{\mathcal{T}}^{n}\phi(x)&=\inf_{\tau}\mathbb{E}\left[\int_{0}^{\tau}e^{-rt}f(X_{t-})\mathrm{d}t+e^{-r\tau}\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{\tau-})\biggl|X_{0-}=x\right].\end{cases}$ (6.3)
The corresponding value function $\tilde{\psi}^{n}$ for the intermediate problem would satisfy the fixed-point problem
$\tilde{\psi}^{n}=\tilde{\mathcal{T}}^{n}\tilde{\psi}^{n}$
as well as the HJB equation
$\min\{\mathcal{L}\tilde{\psi}^{n}-r\tilde{\psi}^{n}+f,\mathcal{M}^{\lambda_{2}}\psi^{n}-\tilde{\psi}^{n}\}=0.$ (6.4)
In fact, one can show by going through the same line of proof as former work (pham2009continuous, , Section 3,5) that for any $n\in\mathbb{N}$, the newly-introduced semi-randomized solution $\tilde{\psi}^{n}$ is the unique viscosity solution to (6.4). Moreover, we have the following propositions.
Proposition 6.4
For any $n\in\mathbb{N}$, $\tilde{\psi}^{n}$ is the unique viscosity solution to (6.4). Moreover, $\tilde{\psi}^{n}(x)\in\mathscr{L}_{\frac{L_{f}}{r-G}}$.
This follows directly from (pham2009continuous, , Section 3,5) and bensoussan2011applications since the obstacle function $\mathcal{M}^{\lambda_{2}}\psi^{n}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$, therefore the proof is omitted.
Proposition 6.5
For any $n\in\mathbb{N}$, $\psi^{n}$, $\tilde{\psi}^{n}$, $\psi^{\bm{\lambda},n}$, $\mathcal{T}^{\bm{\lambda}}\psi^{n}$ are bounded.
We also notice the following property, which bounds the difference between the semi-randomized problem and the classical problem.
Lemma 6.6
For any compact set $K$ with $W\subset K\Subset\mathbb{R}$, for any $n\in\mathbb{N}^{+}$, $\|\psi^{n}-\tilde{\psi}^{n-1}\|_{L^{\infty}(K)}\leq\|\mathcal{M}\psi^{n-1}-\mathcal{M}^{\lambda_{2}}\psi^{n-1}\|_{L^{\infty}(K)}$.
Then we have the following upper bounds for $\mathcal{T}^{\bm{\lambda}}\psi^{n}$ and $\|\mathcal{T}^{\bm{\lambda}}\psi^{n}-\tilde{\psi}^{n}\|_{L^{\infty}(\mathbb{R})}$, respectively.
Proposition 6.7
There exists some constant $C>0$, such that for any $n\in\mathbb{N}$, $x\in\mathbb{R}$, $\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)\leq\mathcal{M}^{\lambda_{2}}\psi^{n}(x)+C\lambda_{1}\log\frac{1}{\lambda_{1}}.$
Proposition 6.8
There exists some constant $C>0$, such that for sufficiently small $\lambda_{1}>0$, for any $n\in\mathbb{N}$, $\|\mathcal{T}^{\bm{\lambda}}\psi^{n}-\tilde{\psi}^{n}\|_{L^{\infty}(\mathbb{R})}\leq C\lambda_{1}\log\frac{1}{\lambda_{1}}$.
Finally, we present the following convergence theorem, which demonstrates that the randomized impulse control problem approximates its classical counterpart as $\bm{\lambda}$ tends to $(0,0)$.
Theorem 6.9
For any compact set $K$ with $W\subset K$, any sequence $\{(\lambda_{1}^{m},\lambda_{2}^{m})\}_{m\geq 1}$ with $\lambda_{1}^{m}\to 0$ and $\lambda_{2}^{m}\to 0$ as $m\to\infty$, there exists a subsequence $\{(\lambda_{1}^{m_{k}},\lambda_{2}^{m_{k}})\}_{k\geq 1}\subset\{(\lambda_{1}^{m},\lambda_{2}^{m})\}_{m\geq 1}$, such that $\|\psi^{\bm{\lambda}^{m_{k}}}(x)-\psi(x)\|_{L^{\infty}(K)}\xrightarrow[]{k\to\infty}0.$
7 Reinforcement Learning for Randomized Impulse Control
Recall that in Section 5, we have provided an iterative approach that eventually converges to the solution to the randomized impulse control problem in Section 5.1 . In this section, we show that this iterative approach actually serves as the foundation for the design of a reinforcement learning algorithm to numerically compute the solution in Definition 2.4 with a temporal–difference (TD) loop. First, we establish a geometric convergence result for the iterative approach under a fixed randomization parameter $\bm{\lambda}\in(0,+\infty)$. Then, we describe the corresponding TD-error algorithm. Finally, we use a linear benchmark problem to illustrate the performances of this algorithm for different $\bm{\lambda}$’s.
7.1 Geometric convergence of the policy improvement algorithm
Theorem 5.2 and 5.3 have guaranteed the convergence for the iterative approach introduced in Section 5.1 to the value function in Definition 2.4. We present the following geometric convergence rate for this approach under the additional standing Assumption 6.1.
Theorem 7.1
There exists some $q\in(0,1)$, such that for any $n\in\mathbb{N}$,
$\|\psi^{\bm{\lambda}}-\psi^{\bm{\lambda},n}\|_{L^{\infty}(\mathbb{R})}\leq\frac{q^{n}}{1-q}\|\psi^{\bm{\lambda}}-\psi^{\bm{\lambda},0}\|_{L^{\infty}(\mathbb{R})}.$
Therefore, the policy improvement scheme (5.1) and (5.2) converges geometrically.
7.2 A Temporal-Difference Error Method
Now, we present a reinforcement learning algorithm for the randomized impulse control problem with a TD loop. In particular, we use a TD scheme to estimate the “policy evaluation” step inside the randomized impulse–control framework.
We fix a uniform step size $\Delta t>0$ and horizon $T=M\Delta t$. At each grid point $t_{i}=i\Delta t$, we propagate
$X_{t_{i+1}}$ $=X_{t_{i}}+b\left(X_{t_{i}}\right)\Delta t+\sigma\left(X_{t_{i}}\right)\sqrt{\Delta t}\,\xi_{i},\quad\xi_{i}\sim\mathcal{N}(0,1),$ (7.1)
$p_{t_{i+1}}$ $=\bigl(1-\pi(X_{t_{i}})\Delta t\bigr)\,p_{t_{i}},$ (7.2)
where (7.1) is the Euler–Maruyama discretization of the controlled SDE and (7.2) tracks the survival probability $p_{t}$, which penalizes premature impulses through the randomized policy $\pi$.
Recall that
$\psi^{\bm{\lambda},n+1}(x)=\mathbb{E}^{x}\left[\int_{0}^{\infty}e^{-rs}p_{s}(f(X_{s})+\pi_{s}\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(X_{s})-\lambda_{1}\mathcal{R}(\pi_{s}))\Big|p_{0}=1\right].$
Assume $\psi^{\bm{\lambda},n+1}$ is approximated via the value network $\psi_{\theta}$ with parameters $\theta$. Then $\psi_{\theta}$ ideally satisfies the one–step Bellman equation
$p_{t_{i}}\psi_{\theta}(X_{t_{i}})=p_{t_{i}}\Bigl[f(X_{t_{i}})+\pi_{t_{i}}\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(X_{t_{i}})-\lambda_{1}\mathcal{R}(\pi_{t_{i}})\Bigr]\Delta t+e^{-r\Delta t}p_{t_{i+1}}\mathbb{E}\bigl[\psi_{\theta}(X_{t_{i+1}})|X_{t_{i}}\bigr].$
Rearranging and using (7.2) yields the instantaneous TD error
$\delta_{i}(\theta)$ $\coloneqq\psi_{\theta}(X_{t_{i}})-e^{-r\Delta t}\bigl(1-\pi_{t_{i}}\Delta t\bigr)\psi_{\theta}(X_{t_{i+1}})-\Bigl[f(X_{t_{i}})+\pi_{t_{i}}\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(X_{t_{i}})-\lambda_{1}\mathcal{R}(\pi_{t_{i}})\Bigr]\Delta t,$ (7.3)
Given $B$ simulated paths $\{X_{t_{i}}^{b}\}_{i=0}^{M}$ ($b=0,1,\dots,B-1)$, the empirical loss minimized via stochastic gradient descent is
[…]
Algorithm 1 summarizes the complete fixed‑point / TD‑error routine. Each outer loop (index $n$) recomputes the Monte‑Carlo estimate of $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}$ and then performs $K$ inner gradient steps with learning rate $\eta$ to minimize (7.4). The next outer loop starts from the last inner iterate, ensuring monotone policy improvement.
Algorithm 1 Reinforcement Learning Algorithm with TD-Error for Randomized Impulse Control
0: randomization parameters $\bm{\lambda}=(\lambda_{1},\lambda_{2})$; neural network $\psi(\cdot;\cdot)$; initial value function $\psi^{0}$; roll-out batch size $B$; time step $\Delta t$ and horizon $T=M\Delta t$; fixed-point iterations NN; gradient steps per outer loop $K$; learning rate $\eta$; initial-state distribution $\rho$ on $x_{\min},x_{\max}]$; SDE drift and diffusion $b$, $\sigma$; entropy function $\mathcal{R}(\cdot)$.
1: Initialize network parameters $\theta_{0}^{0}$ with $\psi(\cdot;\theta_{0}^{0})=\psi^{0}(\cdot)$.
2:for $n=0,1,\dots,N-1$ do
3:  $\psi^{n}(\cdot)\leftarrow\psi(\cdot;\theta_{0}^{n})$.
4:  for $b=0,1,\dots,B-1$ do
5:   Sample $X_{0}^{b}\sim\rho$.
6:   Simulate $\{X_{t_{i}}^{b}\}_{i=0}^{M}$ via the controlled SDE drift/diffusion.
7:   For all $i$, compute $\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t_{i}}^{b})$ (Monte Carlo method).
8:  end for
9:  for $k=0,1,\dots,K-1$ do
10:   for all $(b,i)$ do
11:    $\pi_{t_{i}}^{b}=e^{-\frac{p_{t_{i}}^{b}-\psi(X_{t_{i}}^{b};\theta_{k}^{n})}{\lambda_{1}}}$; clip if needed.
12:   end for
13:   $\mathcal{L}(\theta_{k}^{n})=\frac{1}{B}\sum_{b=0}^{B-1}\sum_{i=0}^{M-1}\Big[\psi(X_{t_{i}}^{b};\theta_{k}^{n})-e^{-r\Delta t}(1-\pi_{t_{i}}^{b}\Delta t)\psi(X_{t_{i+1}}^{b};\theta_{k}^{n})-\Delta t\big(f(X_{t_{i}}^{b})+\pi_{t_{i}}^{b}\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t_{i}}^{b})-\lambda_{1}\mathcal{R}(\pi_{t_{i}}^{b})\big)\Big]^{2}$.
14:   $\theta_{k+1}^{n}\leftarrow\theta_{k}^{n}-\eta\nabla_{\theta}\mathcal{L}(\theta_{k}^{n})$.
15:  end for
16:  $\theta_{0}^{n+1}\leftarrow\theta_{K}^{n}$ .
17:end for
18:
19:return $\psi^{N}(\cdot)=\psi(\cdot;\theta_{0}^{N})$.
7.3 Experimental Setting
We assess Algorithm 1 on the following one–dimensional impulse–control problem, where the drift and volatility of the state dynamic are constants, and the running cost and transaction cost are linear functions.
Specifically, let $b(\cdot)\equiv\mu\in\mathbb{R}$ and $\sigma(\cdot)\equiv\sigma\in\mathbb{R}^{+}$ so that the uncontrolled state dynamic $(X_{t})_{t\geq 0}$ in (2.1) satisfies
$\mathrm{d}X_{t}=\mu\mathrm{d}t+\sigma\mathrm{d}W_{t},\qquad\sigma>0,$
Let $r>0$ be the discount factor. The running cost $f:\mathbb{R}\to\mathbb{R}$ and impulse cost $l:\mathbb{R}\to\mathbb{R}_{+}$ are given by
$f(x)=\begin{cases}hx,&x\geq 0,\\[2.0pt] -px,&x\leq 0,\end{cases}\qquad l(\xi)=\begin{cases}K_{+}+k_{+}\xi,&\xi\geq 0,\\[2.0pt] K_{-}-k_{-}\xi,&\xi<0,\end{cases}$
with constants $h,p,K_{\pm},k_{\pm}>0$. Recall the classical nonlocal operator with test function $\phi$,
$\mathcal{M}\phi(x):=\inf_{\xi\in\mathbb{R}}\{\phi(x+\xi)+l(\xi)\}.$
Finally, assume the nontriviality and growth conditions
$h-rk_{-}>0,\qquad p-rk_{+}>0.$ (7.5)
Here, we specify the following choices of model parameters.
  • Time grid: we fix a total time horizon $T=20$ with step size $\Delta t=0.01$ (i.e., total time steps $M=2000$).
  • Optimization: we adopt a total number of outer fixed‑point iterations $N=30$ with inner gradient steps $K=400$; the optimizer AdamW is set with learning rate $\eta=10^{-3}$ and weight decay $10^{-4}$.
  • Dynamics: constant drift is set to be $\mu=0.03$ and the default volatility is set to be $\sigma=0.2$. The SDE dynamics is sampled via the Euler–Maruyama method:
    $X_{t+\Delta t}-X_{t}=\mu\Delta t+\sigma W_{t,\Delta t},\quad W_{t,\Delta t}\overset{i.i.d.}{\sim}\mathcal{N}(0,\Delta t).$
  • Cost: we use the following cost configuration,
    $f(x)=\begin{cases}x&x\geq 0\\ -x&x<0\end{cases},$ $\ell(x)=\begin{cases}2+0.5x&x\geq 0\\ 2-0.5x&x<0\end{cases},$ discount rate $r=0.1$.
  • Randomization: we vary the randomization parameter $\bm{\lambda}=(\lambda_{1},\lambda_{2})$ within the following choices: $\lambda_{1}\in\{1.0,\,0.5,\,0.1,\,0.05\}$ and $\lambda_{2}=\lambda_{1}$; denote
    $\Lambda\coloneqq\{(1.0,1.0),(0.5,0.5),(0.1,0.1),(0.05,0.05)\}$.
7.4 Experimental Results
We now present the experimental results under the above setting. We are interested in the performance of Algorithm 1 as $\|\bm{\lambda}\|_{\infty}:=\lambda_{1}\vee\lambda_{2}$ decreases and whether the convergence results in Theorem 5.2 and 6.9 can be observed numerically. Besides, we are also interested in the sensitivity analysis of the volatility parameter $\sigma$. This parameter modulates the intrinsic stochasticity of the impulse control problem, regardless of being randomized or not, and we want to demonstrate its impact on the corresponding value function as well as the optimal policy under randomization.
Each figure is the average result out of five runs with distinct random seeds. All simulations are implemented with PyTorch.
7.4.1 Comparison between randomized value functions and their classical counterpart
In Figure 1, we present the comparison results between the randomized and the classical value functions, as $\bm{\lambda}$ varies in $\Lambda$. We plot the corresponding value functions in Figure 1a, where the randomized value functions $\psi^{N}$ are the final outputs of Algorithm 1 and the classical value function $\psi$ is the semi-explicit analytical solution as in constantinides1978existence ; guo2009smooth . It can be observed that as $\bm{\lambda}$ tends to $(0,0)$, the randomized solution approaches the classical benchmark $\psi$, which numerically confirms the theoretical limit $\psi^{\bm{\lambda}}\xrightarrow[]{\|\bm{\lambda}\|_{\infty}\to 0}\psi$ as proved in Theorem 6.9.
For each $\bm{\lambda}\in\Lambda$, we also monitor the evolution of $L_{2}$ relative error between the output of each outer iteration $\psi^{n}$ in Algorithm 1 and the classical value function $\psi$, namely, $\|\psi^{n}-\psi\|_{2}/\|\psi\|_{2}$ for $n=0,\dots,N$. In Figure 1b, we observe that for each $\bm{\lambda}$, this relative error decreases and finally stabilizes as the training proceeds, and such a trend coincides with the convergence of the iterative approach in Theorem 5.2. Moreover, we can see that when $\|\bm{\lambda}\|_{\infty}=0.05$, the lowest among $\Lambda$, the final relative errors reaches the lowest level, which once again is a numerical demonstration of Theorem 6.9; meanwhile, the evolution under $\|\bm{\lambda}\|_{\infty}=0.05$ displays the largest variance. Empirically, it comes from the numerical inaccuracy of the exponential term $\exp(\cdot/\lambda_{1})$ in policy evaluation and $\exp(\cdot/\lambda_{2})$ in Monte-Carlo estimation. Conversely, larger $\|\bm{\lambda}\|_{\infty}$ may introduce bias, but it may enhance stability at the same time. This tradeoff partially explains why $\lambda_{1}=1$ is not the worst performer.
a Value functions.
b $L_{2}$ relative errors evolution in outer iterations.
Figure 1: Comparison in value functions between randomized and classical impulse control with $\bm{\lambda}\in\Lambda$.
7.4.2 Sensitivity Analysis
We investigate how the diffusion volatility $\sigma$ influences the learning dynamics and the resulting randomized solutions. For each $\sigma\in\{0.1,0.2,0.3,0.4\}$, we run Algorithm 1 with $\bm{\lambda}=(0.5,0.5)$ while keeping all other model parameters and hyperparameters fixed, including the training budget, the optimizer, the step size, the number of Monte Carlo samples for the nonlocal term, and the evaluation grid. Results are demonstrated in Figure 2.
For the classical impulse control in constantinides1978existence , the size of the continuation region $|\mathfrak{C}|=O\left(\sigma^{\frac{2}{3}}+\sigma\right)$, and the jump distance $|\xi^{*}(x)|=O\left(\sigma^{\frac{2}{3}}\right)$ for $x$ on the boundary of the continuation region. That is, as $\sigma$ increases, the controller intervenes less frequently, but once hitting the action boundary from inside the continuation region, the jump size is larger. Intuitively, with a larger $\sigma$ value, the controller may make use of the extra randomness when deciding whether to intervene, but may need to compensate for the extra uncertainty by jumping further into the continuation region during each intervention. Consequently, by accumulating more running cost and paying higher cost of intervention, the corresponding value function tends to increase along with $\sigma$. We can observe similar trends under the randomized setting. For instance, in Figure 2a, we observe that the randomized value function monotonically increases as $\sigma$ increases. In Figure 2b, we visualize the optimal intensities $\pi^{*}$ for the survival processes under different values of $\sigma$. As $\sigma$ increases, the flat region where $\pi^{*}$ remains close to 0 (i.e., it is more likely for the controller to wait) keeps expanding. This observation coincides with the enlarging continuation region in the classical problem.
Since $\sigma$ indicates the intrinsic stochasticity of the model, it may require a controller to do more “exploration” as $\sigma$ increases under the randomized setting. In Figure 2b, when $x$ exits the flat region, the intensity of intervention drops as $\sigma$ increases, that is, the controller becomes less committed to exploiting the intervention decision. In Figure 2c, 2d, 2e, we plot the optimal jump distributions at $x\in\{-2,0,2\}$, representing the “jump to the right”, continuation, and “jump to the left” regions in the classical setting, respectively. We can see that, while all the optimal jump distributions $\mu^{*}$ concentrate around the corresponding optimal jumps in the classical setting, as $\sigma$ increases, $\mu^{*}$ displays a larger variance, indicating the tendency to conduct more exploration among possible jumps.
Additional experimental results to compare with the setting $\bm{\lambda}=(1.0,1.0)$ can be found in Appendix B.
a Value functions.
b Intensities of interventions.
c Jumps at x=2.0x=-2.0.
d Jumps at $x=0$.
e Jumps at $x=2.0$.
Figure 2: Sensitivity analysis with respect to volatility $\sigma$ for Algorithm 1 with $\bm{\lambda}=(0.5,0.5)$.
8 Conclusion
In this work, we introduce a novel randomization framework for impulse control problems. Through a compound operator (randomized) consisting of the entropy–regularized versions of the nonlocal and the optimal stopping operators (2.11) and (2.13), we characterized the solution to the randomized problem as a fixed point, which allows us to derive a semi-linear equation (HJB). We rigorously prove the uniqueness of the solution via a verification theorem, thanks to an equivalent randomization scheme with a compound Poisson measure. We also show the existence of the solution using an iterative approach and prove a $\mathcal{C}^{2,\alpha}_{loc}$ regularity for the value function. The existence–uniqueness result thereby ensures the theoretical well-definedness of our randomized problem.
An important finding of this work is the convergence of the randomized solution to its classical counterpart. This, along with the regularity of the value function, shows that our framework is a robust and effective tool for approximating classical impulse control problems. In particular, the iterative approach for the existence proof naturally gives rise to algorithm 1, where the policy improvement steps enjoy a geometric convergence rate. Numerical experiments on a well-studied benchmark problem confirm that the model-free algorithm 1 effectively learns a solution that closely approximates the classical solution. Furthermore, a sensitivity analysis highlights the inherent exploration–exploitation tradeoff within our randomized scheme.
Ultimately, the randomization framework in this work provides not only a sound theoretical basis but also a practical, implementable method for solving complex impulse control problems.
Appendix A Proofs of the Theoretical Results
A.1 Proofs in Section 3
Proof [Proof of Lemma 3.1] Notice that for any $\phi\in\mathscr{F}$, since $\phi$ is bounded from below and is at most of polynomial growth,
$\infty<\inf_{x\in\mathbb{R}}\phi(x)+K\leq\mathcal{M}^{\lambda}\phi(x)\leq\mathbb{E}_{\xi\sim\Phi_{x}}\left[\phi(x+\xi)+l(\xi)\right]<\infty,\quad\forall x\in\mathbb{R}.$
Without loss of generality, assume $\left\|\phi_{1}-\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}<\infty$. For any $x\in\mathbb{R}$ and any $\epsilon>0$, $\exists\mu_{\epsilon}\in\mathcal{P}\left(\mathbb{R}\right)$ such that
$\mathcal{M}^{\lambda}\phi_{2}(x)\leq\mathbb{E}_{\xi\sim\mu_{\epsilon}}\left[\phi_{2}(x+\xi)+l(\xi)\right]+\lambda D_{{\rm KL}}\left(\mu_{\epsilon}\|\Phi_{x}\right)<\mathcal{M}^{\lambda}\phi_{2}(x)+\epsilon,$
and therefore,
$\mathcal{M}^{\lambda}\phi_{1}(x)-\mathcal{M}^{\lambda}\phi_{2}(x)$ $<\mathbb{E}_{\xi\sim\mu_{\epsilon}}\left[\phi_{1}(x+\xi)+l(\xi)\right]-\mathbb{E}_{\xi\sim\mu_{\epsilon}}\left[\phi_{2}(x+\xi)+l(\xi)\right]+\epsilon$
$\leq\left\|\phi_{1}-\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}+\epsilon.$
Similarly, $\mathcal{M}^{\lambda}\phi_{2}(x)-\mathcal{M}^{\lambda}\phi_{1}(x)<\left\|\phi_{1}-\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}+\epsilon$ for any $x\in\mathbb{R}$. Thus,
$\left\|\mathcal{M}^{\lambda}\phi_{1}-\mathcal{M}^{\lambda}\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}\leq\left\|\phi_{1}-\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}+\epsilon,\quad\forall\epsilon>0.$
Sending $\epsilon\to 0$ yields $\left\|\mathcal{M}^{\lambda}\phi_{1}-\mathcal{M}^{\lambda}\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}\leq\left\|\phi_{1}-\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}$. The remaining results follow directly from the definition and particularly the linearity of the operator $\mathcal{M}^{\lambda}$.
Proof [Proof of Lemma 3.2] As $\phi\in\mathscr{F}$, $\mathcal{M}^{\lambda}\phi(x)\in\mathbb{R}$ for all $x\in\mathbb{R}$ by Lemma 3.1-(i). By (2.11),
$\mathcal{M}^{\lambda}\phi(x)=\inf_{\rho\geq 0,\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho(\zeta)\right]}\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho(\zeta)\left(\phi(x+\zeta)+l(\zeta)\right)+\lambda\rho(\zeta)\log\rho(\zeta)\right],\quad\forall x\in\mathbb{R}.$ (A.1)
Notice that the mapping $(0,+\infty)\ni r\mapsto cr+\lambda r\log{r}$ is convex for any constant $c\in\mathbb{R}$. For any $x,\zeta\in\mathbb{R}$, denote
$\tilde{\rho}_{x}(\zeta):=e^{-\frac{[\phi(x+\zeta)-\phi(0)]+[l(\zeta)-l(-x)]}{\lambda}},\quad\hat{\rho}_{x}(\zeta):=e^{-\frac{\phi(x+\zeta)+l(\zeta)}{\lambda}}=e^{-\frac{\phi(0)+l(-x)}{\lambda}}\tilde{\rho}_{x}(\zeta).$
Since $\phi,l\in\mathscr{L}$,
$\tilde{Z}(x)=\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\tilde{\rho}_{x}(\zeta)\right]=\mathbb{E}_{\zeta\sim\mathcal{N}(0,1)}\left[e^{-\frac{[\phi(\zeta)-\phi(0)]+[l(\zeta-x)-l(-x)]}{\lambda}}\right]\in(0,+\infty),\quad\forall x\in\mathbb{R}.$
For any $x\in\mathbb{R}$, with $\hat{Z}(x)=\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\hat{\rho}_{x}(\zeta)\right]=e^{-\frac{\phi(0)+l(-x)}{\lambda}}\tilde{Z}(x)$, we must have
$\mathcal{M}^{\lambda}\phi(x)$ $=\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho_{x}(\zeta)\left(\phi(x+\zeta)+l(\zeta)\right)+\lambda\rho_{x}(\zeta)\log\rho_{x}(\zeta)\right]$
$=-\lambda\log\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\exp\left\{-\frac{\phi(x+\zeta)+l(\zeta)}{\lambda}\right\}\right],$
where $\rho_{x}=\frac{\hat{\rho}_{x}}{\hat{Z}(x)}=\frac{\tilde{\rho}_{x}}{\tilde{Z}(x)}$. Denoting the Lipschitz constant of $\phi$ as $L_{\phi}>0$, for any $x\in\mathbb{R}$ we have
$e^{-\frac{L_{\phi}+L_{l}}{\lambda}|x+\zeta|}\leq\tilde{\rho}_{x}(\zeta)\leq e^{\frac{L_{\phi}+L_{l}}{\lambda}|x+\zeta|},\quad\forall\zeta\in\mathbb{R};$
$\tilde{Z}(x)\geq\underline{Z}\coloneqq\mathbb{E}_{\zeta\sim\mathcal{N}(0,1)}\left[e^{-\frac{L_{\phi}+L_{l}}{\lambda}|\zeta|}\right]>0.$
Take $\mu_{x}\in\mathcal{P}\left(\mathbb{R}\right)$ such that $\mathrm{d}\mu_{x}=\rho_{x}\mathrm{d}\Phi_{x}$. By Markov’s inequality, for any $t\geq 0$, we have
$\mu_{x}\left([-x-t,-x+t]^{\complement}\right)$ $\leq 2\frac{e^{\left(\frac{L_{\phi}+L_{l}}{\sqrt{2}\lambda}\right)^{2}}}{\underline{Z}}\int_{t}^{\infty}\frac{e^{-\frac{1}{2}\left(x-\frac{L_{\phi}+L_{l}}{\lambda}\right)^{2}}}{\sqrt{2\pi}}\mathrm{d}\zeta$
$\leq 2\frac{e^{-\frac{t^{2}}{2}+\frac{L_{\phi}+L_{l}}{\lambda}t}}{\underline{Z}}\leq 2\frac{e^{\frac{1}{2-4\kappa}\left(\frac{L_{\phi}+L_{l}}{\lambda}\right)^{2}}}{\underline{Z}}e^{-\kappa t^{2}},$
for any $\kappa\in\left(0,\frac{1}{2}\right)$. Thus, given $\kappa_{2}\in\left(0,\frac{1}{2}\right)$ and $\kappa_{1}\geq\frac{e^{\frac{1}{2-4\kappa_{2}}\left(\frac{L_{\phi}+L_{l}}{\lambda}\right)^{2}}}{\underline{Z}}$, $\mu_{x}\in{\rm subG}_{x}\left(\kappa_{1},\kappa_{2}\right)$ as in (3.2). Notice that the choices of $(\kappa_{1},\kappa_{2})$ primarily depend on the Lipschitz constants $L_{\phi}$ and $L_{l}$.
Moreover, by (3.1),
$\mathcal{M}^{\lambda}\phi(x)=-\lambda\log{\mathbb{E}_{\zeta\sim\mathcal{N}(0,1)}\left[\exp\left\{-\frac{\phi(\zeta)+l(\zeta-x)}{\lambda}\right\}\right]},\quad x\in\mathbb{R}.$
For any $x,y\in\mathbb{R}$, since $l\in\mathscr{L}_{L_{l}}$,
$l(\zeta-y)\leq l(\zeta-x)+L_{l}|y-x|\text{ and }l(\zeta-x)\leq l(\zeta-y)+L_{l}|x-y|,\quad\forall\zeta\in\mathbb{R}.$
Therefore,
$\mathcal{M}^{\lambda}\phi(y)$ $=-\lambda\log{\mathbb{E}_{\zeta\sim\mathcal{N}(0,1)}\left[\exp\left\{-\frac{\phi(\zeta)+l(\zeta-y)}{\lambda}\right\}\right]}$
$\leq-\lambda\log{\mathbb{E}_{\zeta\sim\mathcal{N}(0,1)}\left[\exp\left\{-\frac{\phi(\zeta)+l(\zeta-x)+L_{l}|y-x|}{\lambda}\right\}\right]}$
$=\mathcal{M}^{\lambda}\phi(x)+L_{l}|y-x|,$
and similarly, $\mathcal{M}^{\lambda}\phi(x)\leq\mathcal{M}^{\lambda}\phi(y)+L_{l}|x-y|$. Thus, $\mathcal{M}^{\lambda}\phi\in\mathscr{L}_{L_{l}}$.
Furthermore, Since $\phi_{1},\phi_{2}\in\mathscr{L}$, there exists $L_{\phi}>0$ such that $\phi_{1},\phi_{2}\in\mathscr{L}_{L_{\phi}}$. Following the arguments above, there exist $\kappa_{1},\kappa_{2}>0$ and $\mu_{1,\cdot},\mu_{2,\cdot}\in{\rm subG}_{\cdot}\left(\kappa_{1},\kappa_{2}\right)$ as in (3.2) such that
$\mathcal{M}^{\lambda}\phi_{i}(x)=\mathbb{E}_{\xi\sim\mu_{i,x}}\left[\phi_{i}(x+\xi)+l(\xi)\right]+\lambda D_{{\rm KL}}\left(\mu_{i,x}\|\Phi_{x}\right),\quad\forall x\in\mathbb{R},\,i\in\{1,2\}.$
Define $\Delta_{R}\coloneqq\left(R+\frac{1}{2\kappa_{2}R}\right)\delta_{R}$ for any $R>0$. Observe that $\lim_{R\to\infty}\Delta_{R}=0$. For any $x\in\mathbb{R}$ and any $\nu\in{\rm subG}_{x}\left(\kappa_{1},\kappa_{2}\right)$,
$\mathbb{E}_{\xi\sim\nu}\left[|x+\xi|\mathbbm{1}\{|x+\xi|>R\}\right]\leq\Delta_{R},\quad\forall R>0.$ (A.2)
Then, for any $x\in\mathbb{R}$ and $R>0$,
$\mathcal{M}^{\lambda}\phi_{1}(x)-\mathcal{M}^{\lambda}\phi_{2}(x)\leq\mathbb{E}_{\xi\sim\mu_{2,x}}\left[\phi_{1}(x+\xi)-\phi_{2}(x+\xi)\right]$
$=$ $\mathbb{E}_{\xi\sim\mu_{2,x}}\left[\left(\mathbbm{1}\left\{|x+\xi|\leq R\right\}+\mathbbm{1}\left\{|x+\xi|>R\right\}\right)\left(\phi_{1}(x+\xi)-\phi_{2}(x+\xi)\right)\right]$
$\leq$ $\mu_{2,x}\left([-R-x,R-x]\right)\|\phi_{1}-\phi_{2}\|_{L^{\infty}([-R,R])}$
$\hskip 15.0pt+\mu_{2,x}\left([-R-x,R-x]^{\complement}\right)|\phi_{1}(0)-\phi_{2}(0)|+2L_{\phi}\mathbb{E}_{\xi\sim\mu_{2,x}}\left[|x+\xi|\mathbbm{1}\left\{|x+\xi|>R\right\}\right]$
$\leq$ $\|\phi_{1}-\phi_{2}\|_{L^{\infty}([-R,R])}+2L_{\phi}\Delta_{R};$
similarly, $\mathcal{M}^{\lambda}\phi_{2}(x)-\mathcal{M}^{\lambda}\phi_{1}(x)\leq\|\phi_{1}-\phi_{2}\|_{L^{\infty}([-R,R])}+2L_{\phi}\Delta_{R}$, and therefore,
$\left\|\mathcal{M}^{\lambda}\phi_{1}-\mathcal{M}^{\lambda}\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}\leq\|\phi_{1}-\phi_{2}\|_{L^{\infty}([-R,R])}+2L_{\phi}\Delta_{R}.$ (A.3)
The conclusion follows with $\mathscr{E}_{R}=2L_{\phi}\Delta_{R}$, $\forall R>0$.
Proof [Proof of Lemma 3.3-(i)] Fix any $0<\lambda<\lambda^{\prime}$. We know that $\exists\mu^{\lambda},\mu^{\lambda^{\prime}}\in\mathcal{P}\left(\mathbb{R}\right)$ such that
$\mathcal{M}^{\gamma}\phi(x)=\mathbb{E}_{\xi\sim\mu^{\gamma}}\left[\phi(x+\xi)+l(\xi)\right]+\gamma D_{{\rm KL}}\left(\mu^{\gamma}\|\Phi_{x}\right),\quad\forall x\in\mathbb{R},$
with $\gamma\in\{\lambda,\lambda^{\prime}\}$. Thus,
$\mathcal{M}^{\lambda}\phi\leq\mathbb{E}_{\xi\sim\mu^{\lambda^{\prime}}}\left[\phi(x+\xi)+l(\xi)\right]+\lambda D_{{\rm KL}}\left(\mu^{\lambda^{\prime}}\|\Phi_{x}\right)\leq\mathcal{M}^{\lambda^{\prime}}\phi(x),\quad\forall x\in\mathbb{R}.$
Proof [Proof of Lemma 3.3-(ii)] For any $y\in\mathbb{R}$, $\delta_{y}\in\mathcal{P}\left(\mathbb{R}\right)$. Thus,
$\mathcal{M}\phi(x)=\inf_{\xi}[\phi(x+\xi)+l(\xi)]=\inf_{\mu\in\mathcal{P}\left(\mathbb{R}\right)}\int_{\mathbb{R}}\phi(x+\xi)+l(\xi)\mu(\mathrm{d}\xi).$
Since $\{\mu\in\mathcal{P}\left(\mathbb{R}\right):\mu\ll\Phi_{x}\}\subset\mathcal{P}\left(\mathbb{R}\right)$ and $D_{{\rm KL}}\left(\mu\|\Phi_{x}\right)\geq 0$ for all $x\in\mathbb{R}$, the conclusion follows.
Proof [Proof of Lemma 3.3-(iii)] By Lemma 3.3-(i) and Lemma 3.3-(ii), it suffices to show that for any $x\in\mathbb{R}$ and any $u>\mathcal{M}\phi(x)$, there exists $\lambda_{0}>0$ such that $\mathcal{M}^{\lambda_{0}}\phi(x)\leq u$. Denote $\mathcal{I}=\{\xi\in\mathbb{R}:\phi(x+\xi)+l(\xi)<u\}$. By continuity of $\phi$ and $l$, $\mathcal{I}\neq\emptyset$ and $\exists[a,b]\in\mathcal{I}$ with $\infty<a<b<+\infty$. Let $\ell=\max_{\xi\in[a,b]}\phi(x+\xi)+l(\xi)$, then $\ell<u$. Denote $c=\frac{a+b}{2}$ and $\delta=b-c$. By Lipschitz condition of $\phi$ and $l$, $\exists\bar{L}>0$ such that $\phi(x+\xi)+l(\xi)<\ell+\bar{L}|\xi-c|,\,\forall\xi\not\in[a,b]$.
Let $\sigma=\frac{u-\ell}{2\bar{L}}\sqrt{\frac{\pi}{2}}>0$ and $\Phi_{\sigma}=\mathcal{N}(c,\sigma)$. Then $\mathbb{E}_{\xi\sim\Phi_{\sigma}}[\phi(x+\xi)+l(\xi)]\leq\ell+L\sigma\sqrt{\frac{2}{\pi}}=\frac{u+\ell}{2}$. Then, for $\lambda_{0}=\frac{u-\ell}{2}\left(-\log{\sigma}+\frac{\sigma^{2}+(c+x)^{2}}{2}\right)^{-1}>0$,
$\mathcal{M}^{\lambda_{0}}\phi(x)\leq\mathbb{E}_{\xi\sim\Phi_{\sigma}}[\phi(x+\xi)+l(\xi)]+\lambda_{0}D_{KL}(\Phi_{\sigma}\|\Phi_{x})\leq u,$
which implies $\lim_{\lambda\to 0^{+}}\mathcal{M}^{\lambda}\phi(x)=\mathcal{M}\phi(x)$ for any $x\in K$.
From (guo2009smooth, , Lemma 2.2-(3)), we know that $\mathcal{M}\phi\in\mathscr{L}$. Then the conclusion follows from Dini’s theorem.
Proof [Proof of Lemma 3.4] Fix any $x,y\in\mathbb{R}$. By Lemma 2.2, for any intensity process $\pi=\{\pi_{t}\}_{t\geq 0}$,
$\left|J^{\lambda,\phi}(x;\pi)-J^{\lambda,\phi}(y;\pi)\right|$
$\leq$ $\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}p_{t}\left(\left|f(X^{x}_{t})-f(X^{y}_{t})\right|+\pi_{t}\left|\phi(X^{x}_{t})-\phi(X^{y}_{t})\right|\right)\mathrm{d}t\right]$
$\leq$ $\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}p_{t}\left(L_{f}+\pi_{t}L_{\phi}\right)\left|X^{x}_{t}-X^{y}_{t}\right|\mathrm{d}t\right]$
$\leq$ $\mathbb{E}\left[\int_{0}^{\infty}e^{-(r-G)t}p_{t}\left(L_{f}+\pi_{t}L_{\phi}\right)\mathrm{d}t\right]|x-y|$
$\leq$ $\max\left\{\frac{L_{f}}{r-G},L_{\phi}\right\}|x-y|\mathbb{E}\left[\int_{0}^{\infty}e^{-(r-G)t}p_{t}\left(r-G+\pi_{t}\right)\mathrm{d}t\right]$
$=$ $\max\left\{\frac{L_{f}}{r-G},L_{\phi}\right\}|x-y|.$
For any $\epsilon>0$, there exists an intensity process $\pi^{y}=\{\pi^{y}_{t}\}_{t\geq 0}$ such that
$T^{\lambda}\phi(y)\leq J^{\lambda,\phi}(y;\pi^{y})<T^{\lambda}\phi(y)+\epsilon.$
Denoting $p^{y}_{t}=\exp\left\{-\int_{0}^{t}\pi^{y}_{s}\mathrm{d}s\right\}$ for all $t\geq 0$, we have
$T^{\lambda}\phi(x)-T^{\lambda}\phi(y)$ $<J^{\lambda,\phi}(x;\pi^{y})-J^{\lambda,\phi}(y;\pi^{y})+\epsilon\leq\max\left\{\frac{L_{f}}{r-G},L_{\phi}\right\}|x-y|+\epsilon;$
Similarly, $T^{\lambda}\phi(y)-T^{\lambda}\phi(x)<\max\left\{\frac{L_{f}}{r-G},L_{\phi}\right\}|x-y|+\epsilon$. Thus,
$|T^{\lambda}\phi(x)-T^{\lambda}\phi(y)|<\max\left\{\frac{L_{f}}{r-G},L_{\phi}\right\}|x-y|+\epsilon,\quad\forall\epsilon>0.$
Sending $\epsilon\to 0$ yields closes the proof.
Theorems A.1 and A.2 that we are about to introduce will immediately imply Theorem 3.5.
Theorem A.1
Suppose $\phi\in\mathscr{L}$ with lower bound $l_{\phi}\geq 0$. There exists a solution $V^{\lambda,\phi}$ to (3.5) such that $V^{\lambda,\phi}\in\mathcal{C}_{loc}^{2,\alpha}(\mathbb{R})$ for any $\alpha\in(0,1)$, and $\lim\limits_{|x|\to\infty}\left|\frac{V^{\lambda,\phi}(x)}{x}\right|<\infty$.
Proof First, we fix any $N,M>0$. Let $\Omega_{N}=(-N,N)$ and $\phi_{M}$ be a smooth cut-off function such that $\phi(x)=e^{x}$ when $x\leq M$ and $\phi_{M}^{x}=e^{M+1}$ when $x>M+1$. Hence, $\phi_{M}$ is bounded and Lipschitz.
Then, we consider the following elliptic equation,
$\begin{cases}(\mathcal{L}-r)v+f-\lambda\phi_{M}\left(-\frac{\phi-v}{\lambda}\right)=0,&x\in\Omega_{N},\\ v(x)=\phi(x),&x\not\in\Omega_{N},\end{cases}$ (A.4)
for $v:\mathbb{R}\to\mathbb{R}$. Let $B(x,y,z)=b(x)z-ry-\lambda\phi_{M}\left(-\frac{\phi(x)-y}{\lambda}\right)+f(x)$ for any $(x,y,z)\in\bar{\Omega}_{N}\times\mathbb{R}\times\mathbb{R}$. We can rewrite (A.4) into
$\frac{\sigma(x)^{2}}{2}v_{N}^{\prime\prime}(x)+B\left(x,v_{N},v_{N}^{\prime}\right)=0.$ (A.5)
Together with the standing Assumptions 2.1 and 2.3, we can apply existence theorems for quasi-linear elliptic equations (see (gilbarg2013elliptic, , Theorem 15.10.) for instance) as well as the comparison principle (see (gilbarg2013elliptic, , Theorem 10.1) for instance) to conclude that (A.4) admits a unique classical solution $v_{N}\in\mathcal{C}_{loc}^{2,\alpha}\cap\mathcal{C}$ for any $\alpha\in(0,1)$. Notice that on $\bar{\Omega}_{N}\times\mathbb{R}\times\mathbb{R}$,
${\rm sign}(y)\lambda\phi_{M}\left(-\frac{\phi(x)-y}{\lambda}\right)\leq\begin{cases}0,&\text{if }y\geq 0;\\ \lambda e^{-\frac{l_{\phi}}{r}},&\text{otherwise},\end{cases}$
and therefore,
$B(x,y,z){\rm sign}(y)=f(x){\rm sign}(y)-r|y|-{\rm sign}(y)\lambda\phi_{M}\left(-\frac{\phi(x)-y}{\lambda}\right)+b(x)|y|.$
Thus,
$\frac{B(x,y,z){\rm sign}(y)}{\frac{\sigma(x)^{2}}{2}z^{2}}\leq\frac{\sigma_{0}^{2}}{2}\frac{|z|\cdot\sup_{x\in\Omega_{N}}b(x)+[\lambda e^{-\frac{l_{\phi}}{\lambda}}+\sup_{x\in\Omega_{N}}f(x)]}{z^{2}}.$
Notice that the constants $C_{1}=\frac{\sigma_{0}^{2}}{2}\cdot\sup_{x\in\Omega_{N}}b(x)$ and $C_{2}=\frac{\sigma_{0}^{2}}{2}[\lambda e^{-\frac{l_{\phi}}{\lambda}}+\sup_{x\in\Omega_{N}}f(x)]\geq 0$ depend only on $(L,\sigma_{0},L_{f},l_{\phi},\lambda,N)$. By maximum principle (gilbarg2013elliptic, , Theorem 10.3), $v_{N}$ is bounded uniformly for all $M>0$. Therefore, given $M$ sufficiently large, $v_{N}$ satisfies (3.5) on $\Omega_{N}$.
Now, choosing any $q>1$ and any $\Gamma\Subset\Omega_{N}$ and applying interior estimate (for example, (gilbarg2013elliptic, , Theorem 9.11.)), we have
$\|v_{N}\|_{\mathcal{W}^{2,q}(\Gamma)}\leq c\left[\|v_{N}\|_{L^{q}(\Gamma)}+\|\lambda e^{-\frac{\phi-v_{N}}{\lambda}}\|_{L^{q}(\Gamma)}\right],$
where $\|v_{N}\|_{L^{q}(\Gamma)}=\left(\int_{\Gamma}|v_{N}(x)|^{q}\mathrm{d}x\right)^{\frac{1}{q}}$, and $\|v_{n}\|_{\mathcal{W}^{2,q}(\Gamma)}=\left(\sum_{k=0}^{2}\|v_{N}^{(k)}\|_{L^{q}(\Gamma)}^{q}\right)^{\frac{1}{q}}$, with some constant $c>0$. By the boundedness of $v_{N}$ on $\Omega_{N}$, we have that
$\|v_{N}\|_{\mathcal{W}^{2,q}(\Gamma)}\leq c^{\prime}\|v_{N}\|_{L^{\infty}(\Gamma)},$
with some constant $c^{\prime}>0$. By Sobolev embedding theorem (see (Evans1998, , Section 5, Theorem 6) for instance), we have that
$\|v_{N}\|_{\mathcal{C}^{1,\alpha}\left(\Gamma_{N}\right)}\leq c^{\prime\prime}\|v_{N}\|_{\mathcal{W}^{2,q}(\Gamma_{N})},$
for any $\alpha\leq 1-\frac{1}{q}$. Thus, $\{v_{N}\}_{N\in\mathbb{N}^{+}}$ is equicontinuous locally. Thus, upon taking subsequences, $\{v_{N}\}_{N\in\mathbb{N}^{+}}$ as a locally uniform limit $V^{\lambda,\phi}$, which is also a weak limit in $\mathcal{W}^{2}_{loc}$. Thus, $V^{\lambda,\phi}\in\mathcal{C}_{loc}^{2,\alpha}$ for any $\alpha\in(0,1)$, since $q>0$ can be arbitrarily large. Finally, it is straightforward to verify that $V^{\lambda,\phi}$ satisfies (3.5) with at most linear growth.
Theorem A.2
Suppose $v\in\mathcal{C}^{2}$ is a classical solution to (3.5) with at most linear growth. Then, $v=T^{\lambda}\phi$.
Proof For any intensity process $\pi=\{\pi_{t}\}_{t\geq 0}$ and the corresponding $p=\{p_{t}=e^{-\int_{0}^{t}\pi_{s}\mathrm{d}s}\}_{t\geq 0}$, applying Itô’s formula to $f(t,X_{t})$ subject to (2.1) with $X_{0}=x\in\mathbb{R}$ with $f(t,y)=e^{-rt}p_{t}v(y)$ for all $t\geq 0$ and $y\in\mathbb{R}$, we have
$e^{-rt}p_{t}v(X_{t})=v(x)+\int_{0}^{t}e^{-rs}p_{s}\left(\mathcal{L}-r-\pi_{s}\right)v(X_{s})\mathrm{d}s+\int_{0}^{t}e^{-rs}p_{s}v^{\prime}(X_{s})\sigma(X_{s})\mathrm{d}W_{s}.$
Since $v$ solves (3.5), and note that $\{\int_{0}^{t}e^{-rs}p_{s}v^{\prime}(X_{s})\sigma(X_{s})\mathrm{d}W_{s}\}_{t\geq 0}$ is a martingale, we have by taking expectation
$v(x)\leq\mathbb{E}_{x}\left[\int_{0}^{t}e^{-rs}p_{s}\left[f(X_{s})-\lambda\mathcal{R}(\pi_{s})+\pi_{s}\phi(X_{s})\right]\mathrm{d}s\right]+\mathbb{E}_{x}\left[e^{-rt}p_{t}v(X_{t})\right].$
By Assumption 2.3-(iii), $\lim_{t\to\infty}\mathbb{E}_{x}\left[e^{-rt}p_{t}v(X_{t})\right]=0$. Therefore, for any $\pi$,
$v(x)\leq J^{\lambda,\phi}(x;\pi),\ \ \forall x\in\mathbb{R}\implies v\leq T^{\lambda}\phi.$
Let $\bar{\pi}=\exp\left\{-\frac{\phi-v}{\lambda}\right\}$ and $\pi^{*}_{t}=\bar{\pi}(X_{t})$ for all $t\geq 0$. Then we have
$v(x)=J^{\lambda,\phi}(x;\pi^{*}),\ \ \forall x\in\mathbb{R}\implies v=T^{\lambda}\phi.$
Proof [Proof of Lemma 3.6-(ii)] Notice that $\mathcal{M}^{\lambda_{2}}\phi\in\mathscr{L}_{L_{l}}$. By Assumption 2.3-(iii), $\frac{L_{f}}{r-G}=\max\left\{\frac{L_{f}}{r-G},L_{l}\right\}$. The conclusion follows from Lemma 3.4.
Proof [Proof of Lemma 3.6-(iii)] Without loss of generality, assume that $\|\phi_{1}-\phi_{2}\|_{L^{\infty}(\mathbb{R})}<\infty$. Fix any $x\in\mathbb{R}$. By Lemma 3.1-(ii), for any intensity process $\pi$,
$\left|J^{\lambda_{1},\mathcal{M}^{\lambda_{2}}\phi_{1}}(x;\pi)-J^{\lambda_{1},\mathcal{M}^{\lambda_{2}}\phi_{1}}(x;\pi)\right|$ $\leq\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-rt}p_{t}\pi_{t}\left|\mathcal{M}^{\lambda_{2}}\phi(X_{t})-\mathcal{M}^{\lambda_{2}}\phi(X_{t})\right|\mathrm{d}t\right]$
$\leq\|\phi_{1}-\phi_{2}\|_{L^{\infty}(\mathbb{R})}\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}p_{t}\pi_{t}\mathrm{d}t\right]$
$\leq\|\phi_{1}-\phi_{2}\|_{L^{\infty}(\mathbb{R})}.$
Similar to the argument in the proof of Lemma 3.4, we have
$|\mathcal{T}^{\bm{\lambda}}\phi_{1}(x)-\mathcal{T}^{\bm{\lambda}}\phi_{2}(x)|\leq\|\phi_{1}-\phi_{2}\|_{L^{\infty}(\mathbb{R})}.$
The conclusion follows.
A.2 Proofs in Section 4
Proof [Proof of Theorem 4.1] We know that $V\in\mathscr{L}$ with a finite lower bound. By Theorem 3.5, $\mathcal{M}^{\lambda_{2}}V\in\mathscr{L}$ with a finite lower bound. Recall that $\mathcal{T}^{\bm{\lambda}}V=T^{\lambda_{1}}[\mathcal{M}^{\lambda_{2}}V]$, hence $\mathcal{T}^{\bm{\lambda}}V$ is a unique classical solution in $\mathcal{C}^{2}\cap\mathscr{L}$ to
$(\mathcal{L}-r)v+f-\lambda_{1}\exp\left\{-\frac{\mathcal{M}^{\lambda_{2}}V-v}{\lambda_{1}}\right\}=0,$ (A.6)
for $v:\mathbb{R}\to\mathbb{R}$. Since $V=\mathcal{T}^{\bm{\lambda}}V$, we must have $V\in\mathcal{C}^{2}$ and it satisfies in a classical sense that
$(\mathcal{L}-r)V+f-\lambda_{1}\exp\left\{-\frac{\mathcal{M}^{\lambda_{2}}V-V}{\lambda_{1}}\right\}=0.$
The conclusion follows.
Proof [Proof of Lemma 4.2] Following the classical dynamical programming principle argument in, for instance, oksendal2007applied , we have the following the HJB equation associated with (4.9), where for any $x\in\mathbb{R}$,
$0$ $=(\mathcal{L}-r)v(x)+f(x)+\inf_{\pi\geq 0;\,\rho\geq 0,\mathbb{E}_{\zeta\sim\Phi_{x},\rho(\zeta)}=1}$ (A.7)
$\hskip 40.0pt\big\{\pi(x)\big[\mathbb{E}_{\zeta\sim\Phi_{x}}\big[\rho(\zeta)\big(v(x+\zeta)+l(\zeta)+\lambda_{2}\log{\rho(\zeta)}\big)-v(x)\big]-\lambda_{1}\mathcal{R}(\pi(x))\big\},$
for $v:\mathbb{R}\to\mathbb{R}$. By linearity in $\pi$, (A.7) is equivalent to
$0$ $=(\mathcal{L}-r)v(x)+f(x)+\inf_{\pi\geq 0}\Bigg\{\pi(x)\cdot$ (A.8)
$\inf_{\rho\geq 0,\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho(\zeta)\right]=1}\left[\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho(\zeta)\left(v(x+\zeta)+l(\zeta)+\lambda_{2}\log{\rho(\zeta)}\right)\right]-v(x)\right]-\lambda_{1}\mathcal{R}(\pi(x))\Bigg\}$
$=(\mathcal{L}-r)v(x)+f(x)+\inf_{\pi\geq 0}\left\{\pi(x)\left[\mathcal{M}^{\lambda_{2}}v(x)-v(x)\right]-\lambda_{1}\mathcal{R}(\pi(x))\right\},$
where the last equality is by (A.1). The equivalence with (HJB) follows from the concavity of $\mathcal{R}$ on $0,+\infty)$.
Proof [Proof of Theorem 4.3] For any $(\pi,\bm{\mu})\in\mathcal{A}$, applying Itô’s formula (e.g., (oksendal2007applied, , Theorem 1.14.)) to $g(t,X_{t})$ subject to (4.6), with $g(t,y)=e^{-rt}V(y)$ for any $(t,y)\in[0,+\infty)\times\mathbb{R}$, we have
$e^{-rt}v(X_{t})$ $=V(x)+\int_{0}^{t}\hat{\mathcal{L}}^{\pi,\nu}[e^{-rs}V(X_{s-})]\mathrm{d}s+\int_{0}^{t}e^{-rs}V^{\prime}(X_{s-})\sigma(X_{s-})\mathrm{d}W_{s}$
$=V(x)+\int_{0}^{t}e^{-rt}\left\{(\mathcal{L}-r)V(X_{s-})+\pi_{s}\mathbb{E}_{\xi\sim\mu_{t}}\left[V(X_{s-}+\xi)-V(X_{s-})\right]\right\}\mathrm{d}s$
$\hskip 60.0pt+\int_{0}^{t}e^{-rs}V^{\prime}(X_{s})\sigma(X_{s-})\mathrm{d}W_{s},\quad\forall t\geq 0.$
Since $V$ is a classical solution to (HJB), by Lemma 4.2, we have that
$e^{-rt}v(X_{t})$ $\geq V(x)+\int_{0}^{t}e^{-rs}V^{\prime}(X_{s})\sigma(X_{s-})\mathrm{d}W_{s}$
$\hskip 10.0pt-\int_{0}^{t}e^{-rs}\left\{f(X_{s-}-\lambda_{1}\mathcal{R}(\pi_{s})+\pi_{s}\left[\mathbb{E}_{\xi\sim\mu_{s}}\left[l(\xi)\right]-\lambda_{2}\mathcal{H}(\mu_{s};X_{s-})\right])\right\}\mathrm{d}s.$
Since $\{\int_{0}^{t}e^{-rs}V^{\prime}(X_{s})\sigma(X_{s-})\mathrm{d}W_{s}\}_{t\geq 0}$, taking expectation on both sides of the inequality, we have
$V(x)$ $\leq\mathbb{E}_{x}\left[\int_{0}^{t}e^{-rs}\left\{f(X_{s-}-\lambda_{1}\mathcal{R}(\pi_{s})+\pi_{s}\left[\mathbb{E}_{\xi\sim\mu_{s}}\left[l(\xi)\right]-\lambda_{2}\mathcal{H}(\mu_{s};X_{s-})\right])\right\}\mathrm{d}s\right]$
$\hskip 50.0pt+\mathbb{E}_{x}\left[e^{-rt}V(X_{t})\right],\quad\forall t\geq 0.$
By Assumption 2.3-(iii), $\lim_{t\to\infty}\mathbb{E}_{x}\left[e^{-rt}V(X_{t})\right]=0$. Therefore, for any $(\pi,\bm{\mu})\in\mathcal{A}$,
$V(x)\leq J^{\bm{\lambda}}(x;\pi,\bm{\mu}),\ \ \forall x\in\mathbb{R}\implies V\leq\hat{\psi}^{\lambda}.$
Since $V\in\mathscr{L}$ with a lower bound, by Theorem 3.5, $\mathcal{M}^{\lambda_{2}}V(x)$ is attained at $\mu_{x}^{*}$ such that $\rho^{*}_{x}=\frac{d\mu^{*}_{x}}{d\Phi_{x}}=\frac{\exp\left\{-\frac{V(x+\xi)+l(\xi)}{\lambda_{2}}\right\}}{\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\exp\left\{-\frac{V(x+\zeta)+l(\zeta)}{\lambda_{2}}\right\}\right]}$. Denote $\bar{\pi}=\exp\left\{-\frac{\mathcal{M}^{\lambda_{2}}V-V}{\lambda_{1}}\right\}$. Let $\pi^{*}_{t}=\bar{\pi}(X_{t-})$, and $\mu^{*}_{t}=\mu^{*}(X_{t-})$ for all $t\geq 0$. By Lemma 4.2 again, we have
$V(x)=J^{\bm{\lambda}}(x;\pi^{*},\bm{\mu}^{*}),\ \ \forall x\in\mathbb{R}\implies V=\hat{\psi}^{\bm{\lambda}}.$
A.3 Proofs in Section 5
Proof [Proof of Lemma 5.1] Fix any $x,x^{\prime}\in\mathbb{R}$. By Lemma 2.2, for $\psi^{\bm{\lambda},0}$, we have
$\left|\psi^{\bm{\lambda},0}(x)-\psi^{\bm{\lambda},0}(x^{\prime})\right|$ $\leq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}|f(X_{t})-f(X_{t}^{\prime})|\mathrm{d}t\right]$
$\leq L_{f}\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}|X_{t}-X_{t}^{\prime}|\mathrm{d}t\right]\leq\frac{L_{f}}{r-G}|x-x^{\prime}|.$
Therefore, $\psi^{\bm{\lambda},0}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Suppose by induction that $\psi^{\bm{\lambda},n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$ for $n\in\mathbb{N}^{+}$. By Theorem 3.5 and Assumption 2.3-(iii), $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Furthermore, one can conclude that $\psi^{\bm{\lambda},n}=T^{\lambda_{1}}[\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}]\in\mathscr{L}_{\frac{L_{f}}{r-G}}$ based on Lemma 3.4, which closes the proof.
Proof [Proof of Theorem 5.2] By definition, for any $x\in\mathbb{R}$,
$\psi^{\bm{\lambda},1}$ $=\inf_{\pi}\mathbb{E}^{x}\left[\int_{0}^{\infty}e^{-rt}p_{t}\left(f(X_{t})-\lambda_{1}\mathcal{R}(\pi_{t})+\pi_{t}\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},0}(X_{t})\right)\mathrm{d}t\right]$
$\leq J^{\lambda_{1},\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},0}}(x;\pi^{0})=\mathbb{E}^{x}\left[\int_{0}^{\infty}e^{-rt}f(X_{t})\mathrm{d}t\right]=\psi^{\bm{\lambda},0}(x).$
Suppose by induction that $\psi^{\bm{\lambda},n}\leq\psi^{\bm{\lambda},n-1}$ holds on $\mathbb{R}$ for $n\in\mathbb{N}^{+}$. By (5.2), for any $x\in\mathbb{R}$,
$\psi^{\bm{\lambda},n+1}(x)$ $=\inf_{\pi}\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}p_{t}\left(f(X_{t})+\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(X_{t})\pi_{t}-\lambda_{1}\mathcal{R}(\pi_{t})\right)\mathrm{d}t\Big|X_{0}=x\right]$
$\leq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}p^{n}_{t}\left(f(X_{t})+\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(X_{t})\pi^{n}_{t}-\lambda_{1}\mathcal{R}(\pi^{n}_{t})\right)\mathrm{d}t\Big|X_{0}=x\right]$
$\leq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}p^{n}_{t}\left(f(X_{t})+\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(X_{t})\pi^{n}_{t}-\lambda_{1}\mathcal{R}(\pi^{n}_{t})\right)\mathrm{d}t\Big|X_{0}=x\right]$
$=\psi^{\bm{\lambda},n}(x),$
where the first inequality comes from the sub-optimality of $\pi^{n}$, and the second inequality is by Lemma 3.1-(iv). The proof by induction is complete.
With a natural lower bound 0, the existence of a pointwise limit $\hat{\psi}^{\bm{\lambda}}\geq 0$ directly follows from monotone convergence theorem. To see the Lipschitz property, we fix any $x,y\in\mathbb{R}$. For any $\epsilon>0$, there exists $N\in\mathbb{N}^{+}$ such that $|\hat{\psi}^{\bm{\lambda}}(x)-{\psi}^{\bm{\lambda},N}(x)|+|\hat{\psi}^{\bm{\lambda}}(y)-{\psi}^{\bm{\lambda},N}(y)|<\epsilon$. By Lemma 5.1, we have
$|\hat{\psi}^{\bm{\lambda}}(x)-\hat{\psi}^{\bm{\lambda}}(y)|$ $\leq|\hat{\psi}^{\bm{\lambda}}(x)-{\psi}^{\bm{\lambda},N}(x)|+|{\psi}^{\bm{\lambda},N}(x)-{\psi}^{\bm{\lambda},N}(y)|+|\hat{\psi}^{\bm{\lambda}}(y)-{\psi}^{\bm{\lambda},N}(y)|$
$<\frac{L_{f}}{r-G}|x-y|+\epsilon.$
Since $\epsilon>0$ and $x,y\in\mathbb{R}$ are arbitrarily fixed, $\hat{\psi}^{\bm{\lambda}}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$.
Moreover, fix any compact set $K\subset\mathbb{R}$. By Dini’s theorem, monotone pointwise convergence of continuous functions to a continuous limit is uniform. Hence,
$\sup_{x\in K}|\psi^{\bm{\lambda},n}(x)-\hat{\psi}^{\bm{\lambda}}(x)|\xrightarrow[]{n\to\infty}0.$
Because $K$ was arbitrary, the convergence is locally uniform on $\mathbb{R}$.
Proof [Proof of Theorem 5.3] For $\hat{\psi}^{\bm{\lambda}}$, 0 is a natural lower bound. From Theorem 5.2, we know that $\hat{\psi}^{\bm{\lambda}}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. By Theorem 3.5, we know that for each $x\in\mathbb{R}$, $\mathcal{M}^{\lambda_{2}}\hat{\psi}^{\bm{\lambda}}$ is attained at and only at $\mu^{*}_{x}$.
By Theorem 3.5, $\mathcal{M}^{\lambda_{2}}\hat{\psi}^{\bm{\lambda}}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$, and the following optimal control problem,
$\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}(x)=\inf_{\pi}\mathbb{E}^{x}\left[\int_{0}^{\infty}e^{-rt}p_{t}\left(f(X_{t})-\lambda_{1}\mathcal{R}(\pi_{t})+\pi_{t}\mathcal{M}^{\lambda_{2}}\hat{\psi}^{\bm{\lambda}}(X_{t})\right)\mathrm{d}t\right],\quad\forall x\in\mathbb{R},$
subject to (2.1) and (2.12) with $p_{0}=1$, is well-posed. In particular, $\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}$ is the unique $\mathcal{C}^{2}$ solution to
$(\mathcal{L}-r)u+f+\inf_{\pi\geq 0}\left\{\pi\left(\mathcal{M}^{\lambda_{2}}\hat{\psi}^{\bm{\lambda}}-u\right)-\lambda_{1}\mathcal{R}(\pi)\right\}=0,$ (A.9)
and the problem admits the optimal control policy $\pi^{*}=\{\pi^{*}_{t}\}_{t\geq 0}$ such that for any $t\geq 0$, $\pi^{*}_{t}=\hat{\pi}(X_{t})$, where
$\hat{\pi}(x)=\exp\left\{-\frac{\mathcal{M}^{\lambda_{2}}\hat{\psi}^{\bm{\lambda}}(x)-\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}(x)}{\lambda_{1}}\right\},\quad\forall x\in\mathbb{R};$
denote $\hat{p}_{t}=\exp\left\{-\int_{0}^{t}\hat{\pi}_{s}ds\right\}$ for all $t\geq 0$. By Theorem 5.2 and (A.3) with $L_{\phi}=\frac{L_{f}}{r-G}$, for any $n\in\mathbb{N}$ and any $R>0$,
$\hat{\psi}^{\bm{\lambda}}(x)-\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}(x)\leq\psi^{\bm{\lambda},n+1}(x)-\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}(x)$
$\leq$ $\;\mathbb{E}^{x}\left[\int_{0}^{\infty}e^{-rt}\hat{p}_{t}\left(f(X_{t})-\lambda_{1}\mathcal{R}(\hat{\pi}_{t})+\hat{\pi}_{t}\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(X_{t})\right)\mathrm{d}t\right]-\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}(x)$
$=$ $\;\mathbb{E}^{x}\left[\int_{0}^{\infty}e^{-rt}\hat{p}_{t}\hat{\pi}_{t}\left(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(X_{t})-\mathcal{M}^{\lambda_{2}}\hat{\psi}^{\bm{\lambda}}(X_{t})\right)\mathrm{d}t\right]$
$\leq$ $\;\left(\|\psi^{\bm{\lambda},n}-\hat{\psi}^{\bm{\lambda}}\|_{L^{\infty}(\Omega_{R})}+\frac{2L_{f}}{r-G}\Delta_{R}\right)\mathbb{E}^{x}\left[\int_{0}^{\infty}e^{-rt}\hat{p}_{t}\hat{\pi}_{t}\mathrm{d}t\right]$
$\leq$ $\;\|\psi^{\bm{\lambda},n}-\hat{\psi}^{\bm{\lambda}}\|_{L^{\infty}(\Omega_{R})}+\frac{2L_{f}}{r-G}\Delta_{R}.$
Fix an arbitrary $\epsilon>0$. By (3.2) and (A.2), there exists $R_{\epsilon}>0$ such that $\Delta_{R_{\epsilon}}<\frac{r-G}{4L_{f}}\epsilon$. By Theorem 5.2, there exists $N_{\epsilon}\in\mathbb{N}$ such that $\|\psi^{\bm{\lambda},n}-\hat{\psi}^{\bm{\lambda}}\|_{L^{\infty}(\Omega_{R_{\epsilon}})}<\frac{\epsilon}{2}$ for all $n>N_{\epsilon}$. Thus,
$\forall\epsilon>0,\,\hat{\psi}^{\bm{\lambda}}(x)-\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}(x)<\epsilon\implies\hat{\psi}^{\bm{\lambda}}(x)\leq\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}(x).$
By Theorem 5.2 and Lemma 3.6-(i), we have that for any $x\in\mathbb{R}$,
$\forall n\in\mathbb{N},\,\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}(x)\leq\mathcal{T}^{\bm{\lambda}}\psi^{\bm{\lambda},n}(x)=\psi^{\bm{\lambda},n+1}(x)\implies\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}(x)\leq\hat{\psi}^{\bm{\lambda}}(x).$
Thus, we must have $\hat{\psi}^{\bm{\lambda}}=\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}$ on $\mathbb{R}$. Replacing $\mathcal{T}^{\bm{\lambda}}\hat{\psi}^{\bm{\lambda}}$ by $\hat{\psi}^{\bm{\lambda}}$ in $\hat{\pi}$, we get $\bar{\pi}$.
A.4 Proofs in Section 6
Proof [Proof of Lemma 6.3] Fix any $x,x^{\prime}\in\mathbb{R}$. By Lemma 5.1, $\psi^{0}\equiv\psi^{\bm{\lambda},0}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$.
Inductively suppose that $\psi^{n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Then by (guo2009smooth, , Lemma 2.2-(3)), $\mathcal{M}\psi^{n-1}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$. Subsequently, using Lemma 2.2, the following holds for any $\tau$:
$|J^{n}(x;\tau)-J^{n}(x^{\prime};\tau)|$
$\leq$ $\mathbb{E}\left[\int_{0}^{\tau}e^{-rs}L_{f}|X_{s}-X^{\prime}_{s}|\mathrm{d}s+e^{-r\tau}L_{\psi^{n-1}}|X_{\tau}-X^{\prime}_{\tau}|\right]$
$\leq$ $\mathbb{E}\left[\sup_{s}e^{-Cs}|X_{s}-X_{s}^{\prime}|\right]\cdot\left(\int_{0}^{\infty}e^{-(r-G)s}L_{f}\mathrm{d}s+e^{-(r-G)\tau}\frac{L_{f}}{r-G}\right)$
$\leq$ $\frac{L_{f}}{r-G}|x-x^{\prime}|.$
By arbitrariness of $\tau$, $\psi^{n}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$.
Proof [Proof of Proposition 6.5] Fix any $x\in\mathbb{R}$. Setting $\tau^{\prime}=\infty$ in the definition of $\psi^{n}$ and $\tilde{\psi}^{n}$, we have
$0$ $\leq\psi^{n}(x)=\inf_{\tau}\mathbb{E}_{x}\left[\int_{0}^{\tau}e^{-rt}f(X_{t})\mathrm{d}t+e^{-r\tau}\mathcal{M}\psi^{n-1}(X_{\tau})\right]$
$\leq\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-rt}f(X_{t})\mathrm{d}t\right]\leq\frac{\|f\|_{L^{\infty}(\mathbb{R})}}{r}\leq\frac{M}{r}.$
Similarly, replacing $\psi^{n}$ by $\tilde{\psi}^{n}$, and $\mathcal{M}\psi^{n-1}$ by $\mathcal{M}^{\lambda_{2}}\psi^{n-1}$ in the above inequality, we have $0\leq\tilde{\psi}^{n}(x)\leq\frac{M}{r}$.
Setting $\pi^{\prime}\equiv 0$, $p^{\prime}\equiv 1$ in the definition of $\psi^{\bm{\lambda},n+1}$ and $\mathcal{T}^{\bm{\lambda}}\psi^{n}$, we have
$0$ $\leq\psi^{\bm{\lambda},n+1}(x)=\mathcal{T}^{\bm{\lambda}}\psi^{\bm{\lambda},n}(x)\leq\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-rt}f(X_{t})\mathrm{d}t\right]\leq\frac{\|f\|_{L^{\infty}(\mathbb{R})}}{r}\leq\frac{M}{r}.$
Proof [Proof of Lemma 6.6] For any $n\in\mathbb{N}^{+}$, $x\in K$, let $\tau^{*}(x)$ be an optimal stopping time for $\psi^{n}(x)$ as defined in (6.1). Then either $\tau^{*}(x)=0$, or $x\in\mathfrak{C}^{n}$ and $\tau^{*}(x)=\inf\{t\geq 0:X_{t}\notin\mathfrak{C}^{n}\}$, hence $X_{\tau^{*}(x)}\in K$. Using $\tau^{*}(x)$ as an admissible (possibly suboptimal) time for $\tilde{\psi}^{\,n-1}(x)$,
$\tilde{\psi}^{n-1}(x)$ $\leq\mathbb{E}_{x}\left[\int_{0}^{\tau^{*}(x)}e^{-rt}f(X_{t})\mathrm{d}t+e^{-r\tau^{*}(x)}\mathcal{M}^{\lambda_{2}}\psi^{n-1}(X_{\tau^{*}(x)})\right]$
$=\psi^{n}(x)+\mathbb{E}_{x}\left[e^{-r\tau^{*}(x)}\left(\mathcal{M}^{\lambda_{2}}\psi^{n-1}(X_{\tau^{*}(x)})-\mathcal{M}\psi^{n-1}(X_{\tau^{*}(x)})\right)\right]$
$\leq\psi^{n}(x)+\left\|\mathcal{M}\psi^{n-1}-\mathcal{M}^{\lambda_{2}}\psi^{n-1}\right\|_{L^{\infty}(K)}.$
Meanwhile, for any $x\in\mathbb{R}$, we know $\psi^{n}\leq\tilde{\psi}^{n-1}$ by substituting the optimal stopping time associated with $\tilde{\psi}^{n-1}$ into the value function definition of $\psi^{n}$, and using $\mathcal{M}\psi^{n}\leq\mathcal{M}^{\lambda_{2}}\psi^{n}$. This closes the proof.
Proof [Proof of Proposition 6.7] Fix $x\in\mathbb{R}$, then With bounded $b,\sigma$ by $M$, Itô’s formula and BDG’s inequality give
$\mathbb{E}_{x}|X_{t}-x|\leq\|b\|_{L^{\infty}(\mathbb{R})}t+\|\sigma\|_{L^{\infty}(\mathbb{R})}\sqrt{t}\leq M(t+\sqrt{t}).$
Subsequently, by $\tfrac{L_{f}}{r-G}$–Lipschitzness of $\mathcal{M}^{\lambda_{2}}\psi^{n}$,
$\bigl|\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t})-\mathcal{M}^{\lambda_{2}}\psi^{n}(x)\bigr|\leq\frac{L_{f}}{r-G}|X_{t}-x|\leq\frac{L_{f}M}{r-G}(t+\sqrt{t}).$ (A.10)
Let us take the constant suboptimal control $\bar{\pi}=\frac{1}{\lambda_{1}^{2}}$, we have
$\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)$ (A.11)
$\leq$ $\;\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})}\left(f(X_{t})+\frac{1}{\lambda_{1}^{2}}\mathcal{M}^{\lambda_{2}}\psi^{n-1}(X_{t})-\lambda_{1}\mathcal{R}\left(\frac{1}{\lambda_{1}^{2}}\right)\right)\mathrm{d}t\right]$
$\leq$ $\;\mathbb{E}_{x}\bigg[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})}\bigg(\frac{1}{\lambda_{1}^{2}}\left(\mathcal{M}^{\lambda_{2}}\psi^{n-1}(x)+\mathcal{M}^{\lambda_{2}}\psi^{n-1}(X_{t})-\mathcal{M}^{\lambda_{2}}\psi^{n-1}(x)\right)+M\bigg)\mathrm{d}t\bigg]$
$\qquad\qquad+\frac{\lambda_{1}\left(1-2\log(\lambda_{1}^{-1})\right)}{1+r\lambda_{1}^{2}}$
$\leq$ $\;\mathcal{M}^{\lambda_{2}}(x)+M\lambda_{1}^{2}+\mathbb{E}_{x}\left[\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})t}\left(\frac{1}{\lambda_{1}^{2}}\left(\mathcal{M}^{\lambda_{2}}\psi^{n-1}(X_{t})-\mathcal{M}^{\lambda_{2}}\psi^{n-1}(x)\right)\right)\mathrm{d}t\right]$
$\qquad\qquad+C\lambda_{1}\log\frac{1}{\lambda_{1}}$
$\leq$ $\;\mathcal{M}^{\lambda_{2}}\psi^{n}(x)+M\lambda_{1}^{2}+\frac{2L_{f}M}{r-G}\lambda_{1}+C\lambda_{1}\log\frac{1}{\lambda_{1}},$
where the second inequality comes from $\|f\|_{L^{\infty}(\mathbb{R})}\leq M$ by Assumption 6.1, the third inequality is due to $\int_{0}^{\infty}e^{-(r+\lambda_{1}^{-2})t}\mathrm{d}t=\frac{\lambda_{1}^{2}}{1+r\lambda_{1}^{2}}\leq C\lambda_{1}^{2}$, and the fourth inequality comes from (A.10). For sufficiently small $\lambda_{1}>0$, we have $\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)\leq\mathcal{M}^{\lambda_{2}}\psi^{n}(x)+C\lambda_{1}\log\frac{1}{\lambda_{1}}.$
Proof [Proof of Proposition 6.8] Recall that $\mathcal{T}^{\bm{\lambda}}\psi^{n}$ and $\tilde{\psi}^{n}$ respectively solves (5.3) and (6.4) with obstacle function $\mathcal{M}^{\lambda_{2}}\psi^{n}$.
Step 1: We prove that $\mathcal{T}^{\bm{\lambda}}\psi^{n}\geq\tilde{\psi}^{n}-C\lambda_{1}$ on $\mathbb{R}$, where $C$ solves $rC=\exp(-C)$.
Let $w=\tilde{\psi}^{n}-C\lambda_{1}$, then
$\mathcal{L}w-rw+f\geq-rC\lambda_{1}=\lambda_{1}\exp(-C)\geq\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{n}-w}{\lambda_{1}}\right),$
where the last inequality is due to $\mathcal{M}^{\lambda_{2}}\psi^{n}-w=\mathcal{M}^{\lambda_{2}}\psi^{n}-\tilde{\psi}^{n}+C\lambda_{1}\geq C\lambda_{1}$ by the second part of (6.4). Applying the maximum principle yields $\mathcal{T}^{\bm{\lambda}}\psi^{n}\geq w$.
Step 2: We prove that $\mathcal{T}^{\bm{\lambda}}\psi^{n}\leq\tilde{\psi}^{n}+C\lambda_{1}\log\frac{1}{\lambda_{1}}$ for some constant $C>0$.
Suppose by contradiction that $\sup_{x}(\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)-\tilde{\psi}^{n}(x)-C\lambda_{1}\log\frac{1}{\lambda_{1}})>0$. Then there exists $x^{\prime}$ such that $\mathcal{T}^{\bm{\lambda}}\psi^{n}(x^{\prime})-\tilde{\psi}^{n}(x^{\prime})-C\lambda_{1}\log\frac{1}{\lambda_{1}}>0$. By choosing a sufficiently small $\epsilon>0$, we have $\mathcal{T}^{\bm{\lambda}}\psi^{n}(x^{\prime})-\tilde{\psi}^{n}(x^{\prime})-C\lambda_{1}\log\frac{1}{\lambda_{1}}-\frac{\epsilon}{2}|x^{\prime}|^{2}>0$. It further holds that $\sup_{x}(\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)-\tilde{\psi}^{n}(x)-C\lambda_{1}\log\frac{1}{\lambda_{1}}-\frac{\epsilon}{2}|x|^{2})>0$. By coercivity, we may assume that the maximum is attained at $x^{*}$.
For simplicity, we set $v(x)=\mathcal{T}^{\bm{\lambda}}\psi^{n}(x)-\tilde{\psi}^{n}(x)-C\lambda_{1}\log\frac{1}{\lambda_{1}}-\frac{\epsilon}{2}|x|^{2}$. By first- and second-order optimality conditions, we have that
$\partial_{xx}v(x^{*})=\partial_{xx}(\mathcal{T}^{\bm{\lambda}}\psi^{n}-\tilde{\psi}^{n})(x^{*})-\epsilon\leq 0;\;\partial_{x}v(x^{*})=\partial_{x}(\mathcal{T}^{\bm{\lambda}}\psi^{n}-\tilde{\psi}^{n})(x^{*})-\epsilon x^{*}=0.$
Hence,
$\mathcal{L}v(x^{*})=\left(b\partial_{x}+\frac{\sigma^{2}}{2}\partial_{xx}\right)v(x^{*})=\mathcal{L}(\mathcal{T}^{\bm{\lambda}}\psi^{n}-\tilde{\psi}^{n})(x^{*})-\frac{\sigma^{2}\epsilon}{2}-b\epsilon x^{*}\leq 0.$
Sending $\epsilon\to 0$ gives $\mathcal{L}(\mathcal{T}^{\bm{\lambda}}\psi^{n}-\tilde{\psi}^{n})(x^{*})\leq 0$. Meanwhile, combining $v(x^{*})>0$ with Proposition 6.7,
$\tilde{\psi}^{n}(x^{*})+C\lambda_{1}\log\frac{1}{\lambda_{1}}+\frac{\epsilon}{2}|x^{*}|^{2}<\mathcal{T}^{\bm{\lambda}}\psi^{n}(x^{*})\leq\mathcal{M}^{\lambda_{2}}\psi^{n}(x^{*})+C_{1}\lambda_{1}\log\frac{1}{\lambda_{1}},$ (A.12)
Then sending $\epsilon\to 0$ yields $\tilde{\psi}^{n}(x^{*})<\mathcal{M}^{\lambda_{2}}\psi^{n}(x^{*})$. Hence, $\mathcal{L}\tilde{\psi}^{n}-r\tilde{\psi}^{n}+f=0$ at $x^{*}$. By (5.3) with $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}$ replaced by $\mathcal{M}^{\lambda_{2}}\psi^{n}$, we have
$\mathcal{L}(\mathcal{T}^{\bm{\lambda}}\psi^{n}-\tilde{\psi}^{n})(x^{*})=r(\mathcal{T}^{\bm{\lambda}}\psi^{n}-\tilde{\psi}^{n})(x^{*})+\lambda_{1}\exp\left(\frac{\mathcal{M}^{\lambda_{2}}\psi^{n}(x^{*})-\mathcal{T}^{\bm{\lambda}}\psi^{n}(x^{*})}{\lambda_{1}}\right).$
By (A.12), the above term is larger than 0, which yields the contradiction.
Proof [Proof of Theorem 6.9] Lower bound: We show that $\psi^{\bm{\lambda},n}\geq\psi^{n}-\frac{\lambda_{1}}{r}$ on $\mathbb{R}$ by induction.
When $n=0$, it holds that $\psi^{\bm{\lambda},0}\equiv\psi^{0}$. Suppose $\psi^{\bm{\lambda},n}\geq\psi^{n}-\frac{\lambda_{1}}{r}$. Set $w^{n}=\psi^{n}-\frac{\lambda_{1}}{r}$. By the value iteration (6.2), we have
$\mathcal{M}^{\lambda_{2}}w^{n}=\mathcal{M}^{\lambda_{2}}\left(\psi^{n}-\frac{\lambda_{1}}{r}\right)=\mathcal{M}^{\lambda_{2}}\psi^{n}-\frac{\lambda_{1}}{r}\geq\psi^{n+1}-\frac{\lambda_{1}}{r}=w^{n+1},$
which implies $\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}w^{n}-w^{n+1}}{\lambda_{1}}\right)\leq\lambda_{1}.$ Combining the above inequality with (6.2), we further have
$\mathcal{L}w^{n+1}-rw^{n+1}+f=\mathcal{L}\psi^{n+1}-r\psi^{n+1}+f+\lambda_{1}\geq\lambda_{1}$
$\geq$ $\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}w^{n}-w^{n+1}}{\lambda_{1}}\right)\geq\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}-w^{n+1}}{\lambda_{1}}\right),$
where the last inequality is due to the induction assumption. Furthermore, recall by (5.3) that
$\mathcal{L}\psi^{\bm{\lambda},n+1}-r\psi^{\bm{\lambda},n+1}+f-\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}-\psi^{\bm{\lambda},n+1}}{\lambda_{1}}\right)=0.$
By comparison principle, we have $\psi^{\bm{\lambda},n+1}\geq w^{n+1}=\psi^{n+1}-\frac{\lambda_{1}}{r}$, which closes the proof.
Returning to the lower bound, let $n\to\infty$ in the claim, we have $\psi^{\bm{\lambda}}\geq\psi-\frac{\lambda_{1}}{r}$ on $\mathbb{R}$.
Upper bound: Fix some $x\in K$. By Proposition 6.2, $\psi^{n}-\psi\leq C(1-\mu)^{n}$ on $\mathbb{R}$. Moreover, the policy iteration gives $\psi^{\bm{\lambda}}\leq\psi^{\bm{\lambda},n}$. Hence,
$\psi^{\bm{\lambda}}(x)-\psi(x)$ $=(\psi^{\bm{\lambda}}(x)-\psi^{\bm{\lambda},n}(x))+(\psi^{\bm{\lambda},n}(x)-\psi^{n}(x))+(\psi^{n}(x)-\psi(x))$ (A.13)
$\leq(\psi^{\bm{\lambda},n}(x)-\psi^{n}(x))+C(1-\mu)^{n}.$
For every $n\in\mathbb{N}$ and every pair $\bm{\lambda}=(\lambda_{1},\lambda_{2})$, we have
$\psi^{\bm{\lambda},n}(x)-\psi^{n}(x)$ (A.14)
$\leq$ $\;\|\mathcal{T}^{\bm{\lambda}}\psi^{\bm{\lambda},n-1}-\mathcal{T}^{\bm{\lambda}}\psi^{n-1}\|_{L^{\infty}(K)}+\|\mathcal{T}^{\bm{\lambda}}\psi^{n-1}-\tilde{\psi}^{n-1}\|_{L^{\infty}(K)}+\|\tilde{\psi}^{n-1}-\psi^{n}\|_{L^{\infty}(K)}$
$\leq$ $\;\|\psi^{\bm{\lambda},n-1}-\psi^{n-1}\|_{L^{\infty}(K)}+C\lambda_{1}\log\frac{1}{\lambda_{1}}+\|\mathcal{M}^{\lambda_{2}}\psi^{n-1}-\mathcal{M}\psi^{n-1}\|_{L^{\infty}(K)},$
where the first inequality comes from triangular inequality, and the second inequality is due to Lemma 3.6-(iii), Proposition 6.8, and Lemma 6.6.
Iteratively applying the arguments of (A.14), and note that $\psi^{\bm{\lambda},0}=\psi^{0}$, we have
$\psi^{\bm{\lambda},n}(x)-\psi^{n}(x)$ (A.15)
$\leq$ $\;\|\psi^{\bm{\lambda},0}-\psi^{0}\|_{L^{\infty}(K)}+nC\lambda_{1}\log\frac{1}{\lambda_{1}}+\sum_{i=1}^{n}\|\mathcal{M}^{\lambda_{2}}\psi^{i-1}-\mathcal{M}\psi^{i-1}\|_{L^{\infty}(K)}$
$=$ $\;nC\lambda_{1}\log\frac{1}{\lambda_{1}}+\sum_{i=1}^{n}\|\mathcal{M}^{\lambda_{2}}\psi^{i-1}-\mathcal{M}\psi^{i-1}\|_{L^{\infty}(K)}.$
For each integer $j\geq 1$ set $\Delta_{j}(\lambda_{2}):=\max_{1\leq i\leq j}\bigl\|\mathcal{M}^{\lambda_{2}}\psi^{i-1}-\mathcal{M}\psi^{i-1}\bigr\|_{L^{\infty}(K)}.$
Lemma 3.3-(iii) implies $\Delta_{j}(\lambda_{2})\to 0$ as $\lambda_{2}\to 0$ for fixed $j$. Choose a decreasing sequence $\eta_{j}\downarrow 0$ such that $\lambda_{2}\leq\eta_{j}$, we have $\Delta_{j}(\lambda_{2})<2^{-j}$. Set $n(\lambda_{1}):=\Bigl\lceil\frac{1}{|\log(1-\mu)|}\log\frac{1}{\lambda_{1}}\Bigr\rceil$, let $\lambda_{2}\leq\eta_{n(\lambda_{1})}$, we then have
$\sum_{i=1}^{n(\lambda_{1})}\bigl\|\mathcal{M}^{\lambda_{2}}\psi^{i-1}-\mathcal{M}\psi^{i-1}\bigr\|_{L^{\infty}(K)}\leq n(\lambda_{1})2^{-n(\lambda_{1})}.$
Subsequently, for sufficiently large $C^{\prime}>0$, one can derive from (A.14) that
$\psi^{\bm{\lambda},n(\lambda_{1})}(x)-\psi^{n(\lambda_{1})}(x)\leq n(\lambda_{1})C\lambda_{1}\log\frac{1}{\lambda_{1}}+n(\lambda_{1})2^{-n(\lambda_{1})}.$
Substituting back to (A.13), we have
$\psi^{\bm{\lambda}}(x)-\psi(x)\leq n(\lambda_{1})C^{\prime}\lambda_{1}\log\frac{1}{\lambda_{1}}+n(\lambda_{1})2^{-n(\lambda_{1})}+C(1-\mu)^{n(\lambda_{1})}.$ (A.16)
Put $\alpha:=\lambda_{1}\log(1/\lambda_{1})$ and let $\widetilde{\mu}:=\max\{1-\mu,\tfrac{1}{2}\}<1$. Both exponential terms on the RHS of (A.16) satisfy $2^{-n(\lambda_{1})},(1-\mu)^{n(\lambda_{1})}\leq\widetilde{\mu}^{\,n(\lambda_{1})}$. Applying (reisinger2020error, , Lemma 4.9) with $\gamma=1$ yields a constant $C_{0}$ such that
$\alpha n(\lambda_{1})+\widetilde{\mu}^{n(\lambda_{1})}\leq C_{0}\alpha\bigl(-\log\alpha\bigr)\quad(\lambda_{1}\to 0).$
Hence the right–hand side of (A.16) is bounded by $C\alpha(-\log\alpha)$, i.e.
$\psi^{\bm{\lambda}}(x)-\psi(x)\leq C\lambda_{1}\log\frac{1}{\lambda_{1}}\log\Bigl(\frac{1}{\lambda_{1}\log(1/\lambda_{1})}\Bigr),$
whenever $\lambda_{2}\leq\eta_{n(\lambda_{1})}$, and the definition of $\eta_{n(\lambda_{1})}$ guarantees that $\lambda_{1}\to 0$, which simultaneously induces a sequence $\lambda_{2}\to 0$. Together with the lower bound $\psi^{\bm{\lambda}}\geq\psi-\lambda_{1}/r$ completes the proof.
A.5 Proof in Section 7
Proof [Proof of Theorem 7.1] Denote $w^{n}(x)=\psi^{\bm{\lambda},n+1}(x)-\psi^{\bm{\lambda},n}(x)$, then for any $x\in\mathbb{R}$,
$\ \mathcal{L}w^{n}(x)-rw^{n}(x)$ $=\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\psi^{\bm{\lambda},n+1}(x)}{\lambda_{1}}\right)-\lambda_{1}\exp\left(-\frac{\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)}{\lambda_{1}}\right)$
$=\exp\left(-\frac{c}{\lambda_{1}}\right)\left(\psi^{\bm{\lambda},n+1}(x)-\psi^{\bm{\lambda},n}(x)-(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x))\right),$
where the second equality holds by mean value theorem, with $c$ lying between $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\psi^{\bm{\lambda},n+1}(x)$ and $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)$.
Note that both $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x)-\psi^{\bm{\lambda},n}(x)$ and $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\psi^{\bm{\lambda},n+1}(x)$ are no more than $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)-\psi^{\bm{\lambda},0}(x)$, we further argue that $\exp(-(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)-\psi^{\bm{\lambda},0}(x))/\lambda_{1})\coloneqq C_{0}$ is bounded, which will be invoked later.
Let us derive the lower bound of $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}$:
$\mathcal{M}^{{\lambda}_{2}}\psi^{\bm{\lambda}}(x)=-\lambda_{2}\log\mathbb{E}_{\xi\sim\Phi_{x}}\!\left[\exp\!\left(-\frac{\psi^{\bm{\lambda}}(x+\xi)+l(\xi)}{\lambda_{2}}\right)\right]$
$=$ $\;\psi^{\bm{\lambda}}(0)+l(-x)-\lambda_{2}\log\mathbb{E}_{\xi\sim\Phi_{x}}\!\left[\exp\!\left(-\frac{\psi^{\bm{\lambda}}(x+\xi)+l(\xi)-\psi^{\bm{\lambda}}(0)-l(-x)}{\lambda_{2}}\right)\right]$
As $\psi^{\bm{\lambda}}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$,
$\exp\!\left(-\frac{\psi^{\bm{\lambda}}(x+\xi)+l(\xi)-\psi^{\bm{\lambda}}(0)-l(-x)}{\lambda_{2}}\right)\leq\exp\left(\frac{\left(\frac{L_{f}}{r-G}+L_{l}\right)|x+\xi|}{\lambda_{2}}\right).$
As $x+\xi\sim\Phi_{0}$, one can derive that its expectation is bounded. This gives the lower boundedness of $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda}}(x)$, which leads to a bound on $C_{0}$. Namely,
$\mathcal{L}w^{n}(x)-rw^{n}(x)\leq C_{0}\left(\psi^{\bm{\lambda},n+1}(x)-\psi^{\bm{\lambda},n}(x)-(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(x)-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(x))\right).$
Since $w^{n}$ is bounded, we may assume that $w^{n}(x)$ attains its maximum at $\overline{x}$, then by first- and second-order optimality conditions
$0$ $\geq\mathcal{L}w^{n}(\overline{x})$
$=\left(r+\exp\left(-\frac{c}{\lambda_{1}}\right)\right)w^{n}(\overline{x})-\exp\left(-\frac{c}{\lambda_{1}}\right)(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(\overline{x})-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(\overline{x})).$
Rearranging the above inequality,
$\max_{x\in\mathbb{R}}w^{n}(x)$ $=w^{n}(\overline{x})\leq\frac{\exp\left(-\frac{c}{\lambda_{1}}\right)}{r+\exp\left(-\frac{c}{\lambda_{1}}\right)}(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(\overline{x})-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(\overline{x}))$ (A.17)
$\leq\left(1-\frac{r}{r+C_{0}}\right)\|\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})}$
$\leq\frac{C_{0}}{r+C_{0}}\|\psi^{\bm{\lambda},n}-\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})},$
where the last inequality is due to the non-expansiveness of $\mathcal{M}^{\lambda_{2}}$ by Lemma 3.1. Following a similar line of proof, we assume that $w^{n}$ attains its minimum at $\underline{x}$, then
$0$ $\leq\mathcal{L}w^{n}(\underline{x})$
$=\left(r+\exp\left(-\frac{c}{\lambda_{1}}\right)\right)w^{n}(\underline{x})-\exp\left(-\frac{c}{\lambda_{1}}\right)(\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n}(\underline{x})-\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}(\underline{x})).$
Thus,
$\min_{x\in\mathbb{R}}w^{n}(x)\leq\frac{C_{0}}{r+C_{0}}\|\psi^{\bm{\lambda},n}-\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})}.$ (A.18)
Combining (A.17) and (A.18) yields
$\|\psi^{\bm{\lambda},n+1}-\psi^{\bm{\lambda},n}\|_{L^{\infty}(\mathbb{R})}\leq\frac{C_{0}}{r+C_{0}}\|\psi^{\bm{\lambda},n}-\psi^{\bm{\lambda},n-1}\|_{L^{\infty}(\mathbb{R})}.$
For simplicity, let $q\coloneqq\frac{C_{0}}{r+C_{0}}$. Fix some $n\in\mathbb{N}^{+}$, for any integer $m>n$, applying the above inequality iteratively, we have
$\|\psi^{\bm{\lambda},m}-\psi^{\bm{\lambda},n}\|_{L^{\infty}(\mathbb{R})}\leq\sum_{k=n}^{m-1}\|\psi^{k+1}-\psi^{k}\|_{L^{\infty}(\mathbb{R})}\leq\frac{q^{n}}{1-q}\|\psi^{\bm{\lambda},1}-\psi^{\bm{\lambda},0}\|_{L^{\infty}(\mathbb{R})}.$
By Theorem 5.2, $\lim_{m\to\infty}\psi^{\bm{\lambda},m}=\psi^{\bm{\lambda}}$, and $\psi^{\bm{\lambda}}\leq\psi^{\bm{\lambda},1}\leq\psi^{\bm{\lambda},0}$. Sending $m\to\infty$, we conclude that for any $n\in\mathbb{N}$,
$\|\psi^{\bm{\lambda}}-\psi^{\bm{\lambda},n}\|_{L^{\infty}(\mathbb{R})}\leq\frac{q^{n}}{1-q}\|\psi^{\bm{\lambda},1}-\psi^{\bm{\lambda},0}\|_{L^{\infty}(\mathbb{R})}\leq\frac{q^{n}}{1-q}\|\psi^{\bm{\lambda}}-\psi^{\bm{\lambda},0}\|_{L^{\infty}(\mathbb{R})}.$
Appendix B Additional Experimental Results
In figs. 3a and 3b, we display the impact of $\sigma$ on the learned value function with respect to two different $\bm{\lambda}$ values. For each $\bm{\lambda}$, the curves retain the same shape and monotone ordering in $\sigma$ (flatter near the origin and higher overall as $\sigma$ increases), reflecting a broader effective continuation region and a reduced marginal value of aggressive impulses. This observation suggests that the impact of $\sigma$ on the randomized value functions enjoys a general consistency across different $\bm{\lambda}$ values. Comparing the two $\bm{\lambda}$ instances, stronger entropy-regularization with $\bm{\lambda}=(1.0,1.0)$ yields a smoother, slightly higher cost profile near the origin, consistent with more conservative interventions; weaker regularization with $\bm{\lambda}=(0.5,0.5)$ permits more aggressive corrections and correspondingly lower minima.
In figs. 3c and 3d, we compare the impact of $\sigma$ on the intervention intensity under the two entropy settings, $\bm{\lambda}=(1.0,1.0)$ (stronger entropy) and $\bm{\lambda}=(0.5,0.5)$ (weaker entropy). Similar the observation for value functions, we also observe the general consistency in the impact of $\sigma$ on the intensity across different values of $\sigma$. For any fixed $\sigma$, the regularization effects under different $\bm{\lambda}$ values are clear. When $\|\bm{\lambda}\|_{\infty}$ is reduced from $1.0$ to $0.5$, we can see that the flat region where the intensity $\pi^{*}$ remains close to 0 visibly enlarges; when the state variable escapes away from the flat region, the increase in the intensity gets more rapid. This suggests that if the regularization effect gets tuned down, the controller will explore less and become more committed to exploiting the continuation/intervention decision.
a Value functions at $\bm{\lambda}=(1.0,1.0)$.
b Value functions at $\bm{\lambda}=(0.5,0.5)$.
c Intensities of interventions at $\bm{\lambda}=(1.0,1.0)$.
d Intensities of interventions at $\bm{\lambda}=(0.5,0.5)$.
Figure 3: Sensitivity analysis with respect to volatility $\sigma$ for Algorithm 1 with $\bm{\lambda}=(1.0,1.0),(0.5,0.5)$.